### v2

# October 17, 2025

Abstract
We provide a complete description of unitarity bounds on the gauge-scalar sectors of models with extra $SU(2)$ doublet, neutral singlet, and charged singlet scalars. Such additions are very frequent in models beyond the Standard Model, and, in particular, they are almost universal in models explaining the dark matter problem. We propose a specific classification and minimal set of scattering matrices containing all the relevant information. We also developed a Mathematica implementation of our results, BounDS, and we use it to fully study a number of simple cases, comparing with the literature, when available. The Mathematica notebook BounDS is provided via a public GitHub repository.
preprint: CFTP/25-xxx
I Introduction
The Standard Model (SM) of particle physics has been extremely successful, culminating in the discovery of a scalar particle consistent with the Higgs boson ATLAS:2012yve ; CMS:2012qbp . Nevertheless, it is not a complete theory, and extensions beyond the Standard Model (BSM) are required. In particular, the SM provides no explanation for dark matter (DM) ParticleDataGroup:2016lqr , whose presence is inferred across a wide range of astronomical scales, from kiloparsecs to the size of the observable Universe. Key probes include the rotation curves of spiral galaxies, the study of galaxy clusters, and measurements of the Cosmic Microwave Background Planck:2018vyg ; Cirelli:2024ssz .
One of the most widely studied scenarios envisions dark matter as consisting of new fundamental particles. The simplest scalar extension introduces a real, neutral $SU(2)$ singlet with a global $\mathbb{Z}_{2}$ symmetry to ensure stability Burgess:2000yq ; GAMBIT:2017gge . Alternatively, a minimal fermionic option is a Weyl fermion, commonly referred to as a sterile neutrino Cirelli:2024ssz . Other proposed models include the Inert Doublet Model, which extends the Higgs sector with a dark doublet Ilnicka:2015jba ; multi-singlet SM extensions Goncalves:2025snm ; the Next-to-Minimal Supersymmetric Standard Model Ellwanger:2009dp ; and models featuring scalar or fermionic portals to hidden sectors Arcadi:2024wwg . Most of these models include extra scalar fields, doublet or singlet under the SM gauge group, and often new discrete symmetries in order to stabilize the dark sector. In exploring such theories, one must start with the theoretical constraints, such as guaranty of a bounded from below potential, assurance of a global minimum, and perturbative unitarity.
In any consistent quantum field theory, probability conservation is encoded in the unitarity of the $S$-matrix Schwartz:2014sze ; Logan:2022uus . In scalar extensions motivated by dark matter, this requirement becomes particularly relevant: new scalar fields can induce scattering amplitudes that grow with energy, and enforcing unitarity imposes bounds on their masses and couplings. For $2\to 2$ scattering processes, unitarity constraints can be applied to the partial-wave decomposition of the amplitudes, with the strongest high-energy limits coming from the zeroth partial-wave amplitude Schwartz:2014sze ; Logan:2022uus . In the high-energy limit, the Equivalence Theorem Veltman:1989ud ensures that the scattering of longitudinally polarized gauge bosons can be replaced by the corresponding Goldstone bosons, allowing these scalar amplitudes to capture the dominant contributions to unitarity bounds. Historically, Lee, Quigg, and Thacker applied this method to derive an upper bound on the Higgs boson mass, $M_{H}^{2}\leq 8\pi\sqrt{2}/(3G_{F})\approx 1\textrm{TeV}^{2}$ Lee:1977yc ; Lee:1977eg . Unitarity has also been applied to the general two-Higgs-doublet model (2HDM) with explicit CP violation Ginzburg:2005dt ; Kanemura:2015ska , to the general $N$-Higgs-doublet model Bento:2017eti , also with fermions Bento:2018fmy , and to constrain large scalar multiplets Hally:2012pu ; Milagre:2024wcg . More recently, perturbative unitarity analyses have been extended beyond the traditional $2\to 2$ limit. For instance, Ref. Bresciani:2025toe studies unitarity in general $M\to N$ scattering processes with $M,N\geq 2$. In particular, in inflationary models featuring a non-minimal coupling with the Ricci scalar, partial-wave unitarity bounds on $2\to N$ processes have improved earlier estimates of the energy scale at which new-physics effective interactions invalidate the perturbative expansion Steingasser:2025txd .
In this work, we extend the analysis of Ref. Bento:2017eti by constructing an $SU(2)\times U(1)$-symmetric model with an arbitrary number of scalar fields (doublets, neutral singlets and charged singlets) and by imposing perturbative unitarity constraints in the high-energy limit. Using partial-wave unitarity for tree-level $2\to 2$ scattering, we derive bounds on the quartic couplings.
In Section II, we present the most general renormalizable $SU(2)\times U(1)$-symmetric model in the high-energy limit, composed of an arbitrary number of $SU(2)$ scalar doublets with hypercharge $\tfrac{1}{2}$ and singlets that are either neutral or carry hypercharge 1. We derive constraints on the Lagrangian couplings imposed by hermiticity and identify additional symmetry relations arising from the interchange of field indices. In Section III, we construct the scattering matrix from second derivatives of the quartic potential, diagonalize it to obtain eigenamplitudes, and impose perturbative unitarity, $|\Lambda|\leq 8\pi$, to constrain the quartic couplings. We propose to classify states by the conserved quantum numbers $|Q,Y,T\rangle$, corresponding to electric charge, hypercharge, and total isospin, respectively. In Appendix A, we illustrate the benefits of labeling states not only by $Q$ and $Y$, but also by total isospin $T$, using the Standard Model as a case study in Appendix B. This classification facilitates the construction of scattering matrices and the identification of independent channels, for which we present a minimal basis (meaning that we discuss the selection of basis states that allows all independent eigenvalues to be determined). We provide explicit formulas for the potential and the matrix elements of all scattering processes allowed in our class of models. In Section IV, we present the Mathematica notebook BounDS that allows for the automatic calculation of the quartic part of the scalar potential and of all scattering matrices in models of this type; the user must only input the particle content, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries [^1] These can either be discrete or continuous, abelian or non-abelian. . We make the notebook publicly available at
https://github.com/andremilagre/BounDS.git
In Section V, we apply these methods to particular models by specifying the scalar content and symmetries, and we compare with results in the literature, when available. We present our results in terms of the parameters of the potential. In specific models, it is sometimes possible to write the parameters of the potential in terms of scalar masses and/or mixing angles. That involves deciding on terms of dimension two and/or three, defining the vacuum of the theory, writing all (neutral and charged) scalar mass matrices, diagonalizing them in order to obtain masses and mixing angles in terms of potential parameters, and then inverting those relations. This must be done on a case-by-case basis, and lies beyond the scope of this work. We present our conclusions in Section VI.
In Appendix C we provide further details on our Mathematica notebook. Finally, Appendix D addresses the inclusion in the basis of additional quantum numbers, when extra flavour symmetries, such as $\mathbb{Z}_{n}$ or CP, are present.
II The Model
II.1 Particle Content
Consider a gauge theory symmetric under the $SU(2)\times U(1)$ group with an arbitrary number of scalar fields (and their charge conjugates) that are either doublets or singlets of $SU(2)$. We define the electric charge $Q$ as the sum of the third component of isospin $T_{3}$, and the hypercharge $Y$, and restrict our discussion to models where the electric charge of the scalar particles is either singly charged [^2] In units of the electron charge $e$. or neutral. As a consequence, this theory may have $n_{D}$ $SU(2)$ scalar doublets with $Y=1/2$ which we denote by
$\Phi_{i}=\begin{pmatrix}\phi_{i}^{+},\,\phi_{i}^{0}\end{pmatrix}^{T},\quad i=1,\ldots,n_{D},$ (1)
$n_{c}$ $SU(2)$ complex scalar singlets with $Y=1$ which we denote by
$\varphi_{i}^{+},\quad i=1,\ldots,n_{c},$ (2)
and $n_{n}$ $SU(2)$ real scalar singlets with $Y=0$ which we denote by
$\chi_{i},\quad i=1,\ldots,n_{n}.$ (3)
Notice that models with $m$ $SU(2)$ complex scalar singlets with $Y=0$ are particular cases of models with $n_{n}=2m$ $SU(2)$ real scalar singlets with $Y=0$; hence, we do not treat them separately. The oblique radiative corrections for this class of models has been thoroughly studied in Refs. Grimus:2007if ; Grimus:2008nb .
II.2 Scalar potential
In this work, we are interested in constraining the parameters of such models by imposing partial wave unitarity bounds on $2\to 2$ scattering processes. We perform all calculations in the high-energy limit where all contributions from propagators vanish, deeming the quartic interactions in the potential the dominant contributions. The most general renormalizable quartic part of the scalar potential of such a theory can be written as: [^3] For the simplest model including the $\kappa_{ab,cd}$ terms, see Sec. V.8 below.
$V\supset V_{4}$ $=$ $\lambda_{ab,cd}\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\left(\Phi_{c}^{\dagger}\Phi_{d}\right)+\alpha_{ab,cd}\left(\varphi^{-}_{a}\varphi^{+}_{b}\right)\left(\varphi^{-}_{c}\varphi^{+}_{d}\right)+\beta_{ab,cd}\left(\chi_{a}\chi_{b}\right)\left(\chi_{c}\chi_{d}\right)+$ (4)
$\delta_{ab,cd}\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\left(\varphi^{-}_{c}\varphi^{+}_{d}\right)+\gamma_{ab,cd}\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\left(\chi_{c}\chi_{d}\right)+\zeta_{ab,cd}\left(\varphi^{-}_{a}\varphi^{+}_{b}\right)\left(\chi_{c}\chi_{d}\right)+$
$\kappa_{ab,cd}\left(\Phi_{a}^{T}\sigma_{2}\Phi_{b}\right)\left(\varphi^{-}_{c}\chi_{d}\right)+\kappa_{ab,cd}^{*}\left(\Phi_{b}^{\dagger}\sigma_{2}\Phi_{a}^{*}\right)\left(\varphi^{+}_{c}\chi_{d}\right),$
where $\sigma_{2}$ is the $2\times 2$ second Pauli matrix, and a sum over repeated indices is implied.
Although compact, not all couplings in this notation are independent. By swapping fields and requiring the scalar potential to be hermitian, we derive the following relations between quartic couplings:
$\lambda_{ab,cd}$ $=$ $\lambda_{ba,dc}^{*}\ =\ \lambda_{cd,ab},$ (5)
$\alpha_{ab,cd}$ $=$ $\alpha_{ba,dc}^{*}\ =\ \alpha_{cd,ab}\ =\ \alpha_{ad,cb},$ (6)
$\beta_{ab,cd}$ $=$ $\beta_{(ab,cd)}^{*}\ =\ \beta_{(ab,cd)},$ (7)
$\delta_{ab,cd}$ $=$ $\delta_{ba,dc}^{*}$ (8)
$\gamma_{ab,cd}$ $=$ $\gamma_{ba,cd}^{*}\ =\ \gamma_{ab,dc},$ (9)
$\zeta_{ab,cd}$ $=$ $\zeta_{ba,cd}^{*}\ =\ \zeta_{ab,dc},$ (10)
$\kappa_{ab,cd}$ $=$ $-\kappa_{ba,cd},$ (11)
where $(ab,cd)$ stands for any permutation of the indices. In particular, these relations further imply
$\lambda_{aa,bb},\ \lambda_{ab,ba}$ $\in$ $\mathbb{R},$ (12)
$\alpha_{aa,bb}=\alpha_{ab,ba}$ $\in$ $\mathbb{R},$ (13)
$\beta_{ab,cd}$ $\in$ $\mathbb{R},$ (14)
$\delta_{aa,bb}$ $\in$ $\mathbb{R},$ (15)
$\gamma_{aa,cd}$ $\in$ $\mathbb{R},$ (16)
$\zeta_{aa,cd}$ $\in$ $\mathbb{R},$ (17)
$\kappa_{aa,cd}$ $=$ $0.$ (18)
III Partial wave unitarity bounds
III.1 Partial wave decomposition
We consider $2\to 2$ scattering processes between the scalars of the theory defined in Section II. Let $A$, $B$, $C$, and $D$ be complex scalar fields and let $a,b,c$, and $d$ be their corresponding flavour indices. At tree-level, the amplitude for the process
$A_{a}\,B_{b}\to C_{c}\,D_{d}$ (19)
may have contributions from $s$, $t$, and $u$-channels, and contact interactions. However, as we take the limit where the Mandelstam variables $s$, $t$, and $u$ go to infinity, the contributions from $s$, $t$, and $u$-channels vanish, respectively  [^4] For an interesting example see, for example, Appendix A of MiguelMSc .. Consequently, in the high-energy limit, only the quartic interactions involving the external scalars contribute to the tree-level amplitude, allowing us to write it as
$\mathcal{M}\left[A_{a}\,B_{b}\to C_{c}\,D_{d}\right]\ =\ -\frac{\partial^{4}V_{4}}{\partial A_{a}\,\partial B_{b}\,\partial C_{c}^{*}\,\partial D_{d}^{*}}.$ (20)
Any square-integrable function may be expressed as an expansion in a complete set of basis functions. To study perturbative unitarity, it is particularly convenient to choose the basis of Legendre polynomials $P_{J}(\cos\theta)$, where $J$ is the total orbital angular momentum of the final state, and $\theta$ the scattering angle. Expanding the amplitude in this basis defines the so-called partial-wave expansion Hally:2012pu ; Milagre:2024wcg :
$\mathcal{M}(\cos\theta)=16\pi\sum_{J=0}^{\infty}a_{J}\left(2J+1\right)P_{J}(\cos\theta).$ (21)
The numerical coefficients $a_{J}$ (also known as partial waves) are defined as
$a_{J}=\frac{2J+1}{32\pi}\int_{-1}^{1}\mathcal{M}(\cos\theta)\,P_{J}(\cos\theta)\ \mathrm{d}\cos\theta,$ (22)
and are constrained by tree-level partial-wave unitarity as Hally:2012pu ; Milagre:2024wcg
[…]
Using the fact that, in the high-energy limit, the amplitude is independent of the scattering angle, and that $P_{0}(\cos\theta)=1$, the most stringent partial-wave unitarity bound arises by applying the second condition of Eq. (23) to the zeroth partial wave $a_{0}$. Thus, using Eq. (20), all quartic couplings must satisfy
$16\pi\left|a_{0}\right|\ =\ \left|N_{ab}N_{cd}\,\frac{\partial^{4}V_{4}}{\partial A_{a}\,\partial B_{b}\,\partial C_{c}^{*}\,\partial D_{d}^{*}}\right|\leq 8\pi,$ (24)
where
$N_{ij}\equiv\frac{1}{\sqrt{2^{\delta_{ij}}}}$ (25)
is a symmetry factor that accounts for identical particles either in the initial or final state.
III.2 Scattering matrices
One may go a step further by employing the method of coupled-channel analysis Logan:2022uus ; Ginzburg:2005dt ; Kanemura:2015ska . This approach takes advantage of the fact that partial wave unitarity bounds may be imposed on any specific process, but also to any superposition of states, provided they have the same quantum numbers. By organizing the zeroth partial waves of such scatterings into a coupled-channel matrix, the unitarity condition translates into a bound on its eigenvalues. The most stringent constraint is obtained by requiring that the modulus of the largest eigenvalue to remain below $8\pi$. Once again, note that $J=0$ partial waves should be multiplied by a $1/\sqrt{2}$ for every pair of identical particles in the initial or final state.
We can classify two-particle initial and final states according to their total electric charge $Q$ and hypercharge $Y$, following the approach in Bento:2017eti . In addition to $Q$ and $Y$, any scattering process involving $SU(2)$ doublets must also conserve total isospin, $T$. Following the results of Appendix A, we list in Table 1 the most minimal set of independent two-particle states, labeled by the quantum numbers $|Q,Y,T\rangle$. Throughout we use the notation
$\phi^{+}_{[i}\,\phi^{0}_{j]}$ $\equiv$ $\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{0}_{j}-\phi^{+}_{j}\,\phi^{0}_{i}\right),$ (26)
$\Phi_{i}\,\Phi_{j}^{*}$ $\equiv$ $\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{-}_{j}+\phi^{0}_{i}\,\phi^{0*}_{j}\right).$ (27)
$|Q,Y,T\rangle$ State Conditions Dimensionality
$|2,2,0\rangle$ $\varphi^{+}_{i}\varphi^{+}_{j}$ $i\leq j$ $n_{c}(n_{c}+1)/2$
$|2,\frac{3}{2},\frac{1}{2}\rangle$ $\phi^{+}_{i}\varphi^{+}_{j}$ $n_{c}n_{D}$
$|2,1,1\rangle$ $\phi^{+}_{i}\phi^{+}_{j}$ $i\leq j$ $n_{D}(n_{D}+1)/2$
$|1,1,0\rangle$ $\left\{\phi^{+}_{[i}\,\phi^{0}_{j]},\ \varphi^{+}_{i}\chi_{j}\right\}$ $\{i<j,\,\textrm{---}\,\}$ $n_{D}(n_{D}-1)/2+n_{c}n_{n}$
$|1,\frac{1}{2},\frac{1}{2}\rangle$ $\left\{\phi^{+}_{i}\chi_{j},\ \phi^{0*}_{i}\varphi^{+}_{j}\right\}$ $n_{D}(n_{c}+n_{n})$
$|1,0,1\rangle$ $\phi^{+}_{i}\phi^{0*}_{j}$ $n_{D}^{2}$
$|0,0,0\rangle$ $\left\{\Phi_{i}\Phi_{j}^{*},\ \varphi_{i}^{+}\varphi_{j}^{-},\ \chi_{i}\chi_{j}\right\}$ $\{\,\textrm{---}\,,\,\textrm{---}\,,\,i\leq j\}$ $n_{n}(n_{n}+1)/2+n_{D}^{2}+n_{c}^{2}$
Table 1: Basis of two-particle states labeled by $|Q,Y,T\rangle$. This table had been reduced by including only non-redundant sets of states; see Appendix A for more details.
Looking at Table 1, we notice that the non-redundant sets of states could be classified exclusively in terms of $Y$ and $T$, as is done for the 2HDM in Ginzburg:2005dt . On the other hand, Ref. Bento:2017eti advocated $Q$ and $Y$ when considering all states, since then, as can be seen in Table 9 of Appendix A, there are states with different $Q$ for the same $(Y,T)$. We find it best to make the bridge and consider $|Q,Y,T\rangle$.
Using Eqs. (4)–(18), and (24), the scattering matrix elements for the relevant processes read
$16\pi\,a_{0}\left[\varphi^{+}_{a}\varphi^{+}_{b}\to\varphi^{+}_{c}\varphi^{+}_{d}\right$] $=$ $4\,N_{ab}N_{cd}\,\alpha_{ca,db},$ (28)
$16\pi\,a_{0}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right$] $=$ $\delta_{ca,db},$ (29)
$16\pi\,a_{0}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right$] $=$ $N_{ab}N_{cd}\,\left(2\lambda_{ca,db}+2\lambda_{da,cb}\right),$ (30)
$16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\phi^{+}_{[c}\,\phi^{0}_{d]}\right$] $=$ $\left(2\lambda_{ca,db}-2\lambda_{da,cb}\right),$ (31)
$16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\varphi^{+}_{c}\chi_{d}\right]$ $=$ $2\sqrt{2}\,i\,\kappa_{ba,cd},$ (32)
$16\pi\,a_{0}\left[\varphi^{+}_{a}\chi_{b}\to\varphi^{+}_{c}\chi_{d}\right]$ $=$ $2\zeta_{ca,bd},$ (33)
$16\pi\,a_{0}\left[\phi^{+}_{a}\chi_{b}\to\phi^{+}_{c}\chi_{d}\right]$ $=$ $2\gamma_{ca,bd},$ (34)
$16\pi\,a_{0}\left[\phi^{+}_{a}\chi_{b}\to\phi^{0*}_{c}\varphi^{+}_{d}\right$] $=$ $2i\kappa_{ca,db},$ (35)
$16\pi\,a_{0}\left[\phi^{0*}_{a}\varphi^{+}_{b}\to\phi^{0*}_{c}\varphi^{+}_{d}\right$] $=$ $\delta_{ac,db},$ (36)
$16\pi\,a_{0}\left[\phi^{+}_{a}\phi^{0*}_{b}\to\phi^{+}_{c}\phi^{0*}_{d}\right]$ $=$ $2\lambda_{ca,bd},$ (37)
$16\pi\,a_{0}\left[\Phi_{a}\,\Phi_{b}^{*}\to\Phi_{c}\,\Phi_{d}^{*}\right]$ $=$ $4\lambda_{ba,cd}+2\lambda_{ca,bd},$ (38)
$16\pi\,a_{0}\left[\Phi_{a}\,\Phi_{b}^{*}\to\varphi^{+}_{c}\varphi^{-}_{d}\right]$ $=$ $\sqrt{2}\delta_{ba,cd},$ (39)
$16\pi\,a_{0}\left[\Phi_{a}\,\Phi_{b}^{*}\to\chi_{c}\,\chi_{d}\right]$ $=$ $2\sqrt{2}N_{cd}\,\gamma_{ba,cd},$ (40)
$16\pi\,a_{0}\left[\varphi^{+}_{a}\varphi^{-}_{b}\to\varphi^{+}_{c}\varphi^{-}_{d}\right]$ $=$ $4\alpha_{ba,cd},$ (41)
$16\pi\,a_{0}\left[\varphi^{+}_{a}\varphi^{-}_{b}\to\chi_{c}\,\chi_{d}\right]$ $=$ $2\,N_{cd}\,\zeta_{ba,cd},$ (42)
$16\pi\,a_{0}\left[\chi_{a}\,\chi_{b}\to\chi_{c}\,\chi_{d}\right]$ $=$ $24\,N_{ab}N_{cd}\,\beta_{ab,cd}.$ (43)
Below, we study specific models, showing the scattering matrices $M_{|Q,Y,T\rangle}$ and, when an analytical expression is possible, the corresponding eigenvalues. Note that in Ref. Bento:2022vsb a method is proposed based on principal minors, that forgoes diagonalization. In specific numerical simulations, this is likely to be computationally preferable in all models dealing with large scattering matrices.
IV Mathematica notebook
Deriving partial-wave unitarity bounds for different models is a rather repetitive task. As stated, for a model specified by definite values of $n_{D}$, $n_{c}$, and $n_{n}$, the procedure requires assembling a scattering matrix for each state $|Q,Y,T\rangle$ listed in Table 1, computing the corresponding eigenvalues, and verifying that they remain below $8\pi$. In models with additional flavour symmetries, all quartic terms forbidden by those symmetries must also be set to zero.
To optimize this process and reduce the burden on the high-energy physics phenomenology community, we have developed the Mathematica tool BounDS that automates the calculation. The user simply has to specify the values of $n_{D}$, $n_{c}$, and $n_{n}$, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries. BounDS carries out the necessary steps to:
  • Compute the set of all linearly independent quartic couplings allowed by the symmetries.
  • Calculate the 7 independent scattering matrices $M_{|Q,Y,T\rangle}$.
  • Block-diagonalize the scattering matrices by swapping rows and columns.
  • Output the quartic part of the scalar potential and scattering matrices in form.
  • Output closed expressions for the eigenvalues of the scattering matrices, when possible.
Additional details can be found in Appendix C and the notebook BounDS can be downloaded from:
https://github.com/andremilagre/BounDS.git
The results of the next section are derived from this notebook.
V Unitarity Bounds for particular cases
We now proceed to study unitarity bounds in specific scenarios by specifying the values of $n_{D}$, $n_{c}$, and $n_{n}$, as well as any additional symmetries. The cases studied explicitly in this work are summarized in Table 2.
$_{D} nn$$_{n}$ $n_{c}$ Symmetries Ref.
1 0 0 Lee:1977eg
1 1 0
2 0 0 $\mathbb{Z}_{2}$ Ginzburg:2005dt
2 0 0 Ginzburg:2005dt
1 2 0 Muhlleitner:2020wwk generalized
2 1 0 $\mathbb{Z}_{2}$
2 2 0 $\mathbb{Z}_{2}\otimes\mathbb{Z}^{\prime}_{2}$ Boto:2025mmn
2 1 1
3 0 0 $\mathbb{Z}_{3}$ Bento:2017eti
Table 2: Models presented as examples in this article.
V.1 The Standard Model
In the Standard Model, the scalar sector consists of a single $SU(2)$ doublet $\Phi_{1}$. Consequently, the quartic part of the scalar potential simply reads
$V_{4}=\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}.$ (44)
V.1.1 Scattering Matrices
For this minimal scalar content, there are only three non-zero scattering matrices. These matrices have rank-1 and take the following form:
$M_{|1,0,1\rangle}=M_{|2,1,1\rangle}=2\lambda_{11,11},$ (45)
$M_{|0,0,0\rangle}=6\lambda_{11,11}.$ (46)
V.1.2 Unitarity Bounds
Applying the partial wave unitarity condition of Eq. (24) to the eigenvalues of these scattering matrices, we find
$|2\lambda_{11,11}|\leq 8\pi,\quad|6\lambda_{11,11}|\leq 8\pi\quad\Longrightarrow\quad\lambda_{11,11}\leq\frac{4\pi}{3}.$ (47)
In order to facilitate the comparison of our results with those of Ref. Bento:2017eti , it is useful to provide a short dictionary of notations. Table 3 lists the correspondence between the coupling employed in this work and the one used in Ref. Bento:2017eti .
Term Our Notation Notation in Bento:2017eti
$(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\lambda$
Table 3: Comparison of coupling notation.
This recovers the classic results of Lee:1977eg .
V.2 1 Scalar Doublet and 1 Neutral Scalar Singlet
We now consider an extension of the SM where, besides the $SU(2)$ scalar doublet $\Phi_{1}$, we add one neutral scalar singlet $\chi_{1}$. In this case, the quartic part of the scalar potential takes the form
$V_{4}=\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\beta_{11,11}\,\chi_{1}^{4}+\gamma_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{1}^{2}.$ (48)
V.2.1 Scattering Matrices
The set of non-zero scattering matrices for this model are given by:
$M_{|1,0,1\rangle}=M_{|2,1,1\rangle}=2\lambda_{11,11},$ (49)
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=2\gamma_{11,11},$ (50)
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&2\sqrt{2}\gamma_{11,11}\\ 2\gamma_{11,11}&12\beta_{11,11}\\ \end{bmatrix}$. (51)
V.2.2 Unitarity Bounds
We apply the partial wave unitarity condition of Eq. (24) to each of the eigenvalues of the zero partial wave amplitude matrix and find
$\left|6\beta_{11,11}+3\lambda_{11,11}\pm\sqrt{9\left(\lambda_{11,11}-2\beta_{11,11}\right){}^{2}+4\gamma_{11,11}^{2}}\right|$ $\leq$ $8\pi,$ (52)
$|2\lambda_{11,11}|$ $\leq$ $8\pi,$ (53)
$|2\gamma_{11,11}|$ $\leq$ $8\pi.$ (54)
V.3 The $\mathbb{Z}_{2}$-Symmetric 2HDM
We now consider a 2HDM model with a discrete $\mathbb{Z}_{2}$ symmetry that acts on the two scalar $SU(2)$ doublets, $\Phi_{1}$ and $\Phi_{2}$, as
$\Phi_{1}\to\Phi_{1},\qquad\Phi_{2}\to-\Phi_{2}.$ (55)
This case illustrates the use of our Mathematica program in the presence of symmetries. Under this symmetry, the quartic part of the scalar potential reads
$V_{4}^{\mathbb{Z}_{2}}$ $=$ $\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+2\lambda_{12,21}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)$ (57)
$\ +\ \lambda_{12,12}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{12,12}^{*}\left(\Phi_{2}^{\dagger}\Phi_{1}\right)^{2}$
$=$ $\frac{\lambda_{1}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\frac{\lambda_{2}}{2}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{3}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+\lambda_{4}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)+\frac{\lambda_{5}}{2}\left[\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\text{h.c.}\right],$ (58)
where, in Eq. (58) we have employed the standard notation of Ref. Ginzburg:2005dt and, without loss of generality, choose $\lambda_{5}$ to be real. The mapping between the two notations is given in Table 4.
Term Our Notation Notation in Ginzburg:2005dt
$(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\frac{\lambda_{1}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{2})^{2}$ $\lambda_{22,22}$ $\frac{\lambda_{2}}{2}$
$(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})$ $2\lambda_{11,22}$ $\lambda_{3}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})$ $2\lambda_{12,21}$ $\lambda_{4}$
$(\Phi_{1}^{\dagger}\Phi_{2})^{2}$ $\lambda_{12,12}$ $\frac{\lambda_{5}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{1})^{2}$ $\lambda_{12,12}^{*}$ $\frac{\lambda_{5}}{2}$
Table 4: Comparison of coupling notation.
V.3.1 Scattering Matrices
The set of non-zero scattering matrices for this model reads
$M_{|2,1,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{5}&0\\ \lambda_{5}&\lambda_{2}&0\\ 0&0&\lambda_{3}+\lambda_{4}\end{bmatrix}$, (59)
$M_{|1,0,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{4}&0&0\\ \lambda_{4}&\lambda_{2}&0&0\\ 0&0&\lambda_{3}&\lambda_{5}\\ 0&0&\lambda_{5}&\lambda_{3}\end{bmatrix}$, (60)
$M_{|0,0,0\rangle}=\begin{bmatrix}3\lambda_{1}&2\lambda_{3}+\lambda_{4}&0&0\\ 2\lambda_{3}+\lambda_{4}&3\lambda_{2}&0&0\\ 0&0&\lambda_{3}+2\lambda_{4}&3\lambda_{5}\\ 0&0&3\lambda_{5}&\lambda_{3}+2\lambda_{4}\end{bmatrix}$, (61)
$M_{|1,1,0\rangle}=\lambda_{3}-\lambda_{4}.$ (62)
V.3.2 Eigenvalues
By imposing the unitarity condition of Eq. (24) to each of the eigenvalues of the zero partial wave amplitude matrix we derive
$\left|\lambda_{3}\pm\lambda_{4}\right|$ $\leq$ $8\pi,$ (63)
$\left|\lambda_{3}\pm\lambda_{5}\right|$ $\leq$ $8\pi,$ (64)
$\left|\lambda_{3}+2\lambda_{4}\pm 3\lambda_{5}\right|$ $\leq$ $8\pi,$ (65)
$\frac{1}{2}\left|\lambda_{1}+\lambda_{2}\pm\sqrt{\left(\lambda_{1}-\lambda_{2}\right)^{2}+4\lambda_{4}^{2}}\right|$ $\leq$ $8\pi,$ (66)
$\frac{1}{2}\left|\lambda_{1}+\lambda_{2}\pm\sqrt{\left(\lambda_{1}-\lambda_{2}\right)^{2}+4\lambda_{5}^{2}}\right|$ $\leq$ $8\pi,$ (67)
$\frac{1}{2}\left|3\lambda_{1}+3\lambda_{2}\pm\sqrt{9\left(\lambda_{1}-\lambda_{2}\right)^{2}+4\left(2\lambda_{3}+\lambda_{4}\right)^{2}}\right|$ $\leq$ $8\pi.$ (68)
We find that our results are in agreement with those derived in Ginzburg:2005dt .
V.4 The general 2HDM
In this case, we consider the most general scalar potential involving two $SU(2)$ doublets, $\Phi_{1}$ and $\Phi_{2}$. The quartic part of the potential is given by:
$V_{4}$ $=$ $\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+2\lambda_{12,21}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)+$ (70)
$\left[\lambda_{12,12}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,12}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{1}^{\dagger}\Phi_{2}\right)+2\lambda_{12,22}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+\text{h.c.}\right]$
$=$ $\frac{\lambda_{1}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\frac{\lambda_{2}}{2}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{3}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+\lambda_{4}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)+$
$\left[\frac{\lambda_{5}}{2}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{6}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{1}^{\dagger}\Phi_{2}\right)+\lambda_{7}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+\text{h.c.}\right],$
where, in Eq. (70) we have employed, once again, the standard notation of Ref. Ginzburg:2005dt . The mapping between the two notations is given in Table 5.
Term Our Notation Notation in Ginzburg:2005dt
$(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\frac{\lambda_{1}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{2})^{2}$ $\lambda_{22,22}$ $\frac{\lambda_{2}}{2}$
$(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})$ $2\lambda_{11,22}$ $\lambda_{3}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})$ $2\lambda_{12,21}$ $\lambda_{4}$
$(\Phi_{1}^{\dagger}\Phi_{2})^{2}$ $\lambda_{12,12}$ $\frac{\lambda_{5}}{2}$
$(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{1}^{\dagger}\Phi_{2})$ $2\lambda_{11,12}$ $\lambda_{6}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{2})$ $2\lambda_{12,22}$ $\lambda_{7}$
Table 5: Comparison of coupling notation.
V.4.1 Scattering Matrices
The set of non-zero scattering matrices for this model reads
$M_{|2,1,1\rangle}=\begin{bmatrix}\lambda_{1}&\sqrt{2}\,\lambda_{6}^{*}&\lambda_{5}^{*}\\ \sqrt{2}\,\lambda_{6}&\lambda_{3}+\lambda_{4}&\sqrt{2}\,\lambda_{7}^{*}\\ \lambda_{5}&\sqrt{2}\,\lambda_{7}&\lambda_{2}\end{bmatrix}$, (71)
$M_{|1,0,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{6}&\lambda_{6}^{*}&\lambda_{4}\\ \lambda_{6}^{*}&\lambda_{3}&\lambda_{5}^{*}&\lambda_{7}^{*}\\ \lambda_{6}&\lambda_{5}&\lambda_{3}&\lambda_{7}\\ \lambda_{4}&\lambda_{7}&\lambda_{7}^{*}&\lambda_{2}\end{bmatrix}$, (72)
$M_{|0,0,0\rangle}=\begin{bmatrix}3\lambda_{1}&3\lambda_{6}^{*}&3\lambda_{6}&2\lambda_{3}+\lambda_{4}\\ 3\lambda_{6}&\lambda_{3}+2\lambda_{4}&3\lambda_{5}&3\lambda_{7}\\ 3\lambda_{6}^{*}&3\lambda_{5}^{*}&\lambda_{3}+2\lambda_{4}&3\lambda_{7}^{*}\\ 2\lambda_{3}+\lambda_{4}&3\lambda_{7}^{*}&3\lambda_{7}&3\lambda_{2}\end{bmatrix}$, (73)
$M_{|1,1,0\rangle}=\lambda_{3}-\lambda_{4}.$ (74)
Our scattering matrices agree with those in Ginzburg:2005dt . If we set $\lambda_{6}=\lambda_{7}=0$ (along with their complex conjugates), we re-obtain, as expected, the results presented in Section V.3.1.
V.4.2 Unitarity Bounds
The eigenvalues of the scattering matrices in this case are, in general, too complex to write down in closed form. However, some can be computed analytically, and unitarity bounds are imposed accordingly,
$|\lambda_{3}-\lambda_{4}|\leq 8\pi.$ (75)
V.5 1 Scalar Doublet and 2 Neutral Scalar Singlets
In this case, the most general quartic part of the scalar potential reads:
$V_{4}=\$$\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\beta_{11,11}\,\chi_{1}^{4}+\beta_{22,22}\,\chi_{2}^{4}+6\,\beta_{11,22}\,\chi_{1}^{2}\chi_{2}^{2}+4\,\beta_{11,12}\,\chi_{1}^{3}\chi_{2}+4\,\beta_{12,22}\,\chi_{1}\chi_{2}^{3}$
$+\gamma_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{1}^{2}+\gamma_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{2}^{2}+2\,\gamma_{11,12}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{1}\chi_{2},$ (76)
where $\Phi_{1}$ denotes the scalar doublet and $\chi_{1}$ and $\chi_{2}$ are the real scalar singlets. We have used the relations in Eqs. (7) and (9) to simplify the quartic part of the potential.
V.5.1 Scattering Matrices
The set of non-zero scattering matrices for this model reads
$M_{|1,0,1\rangle}=M_{|2,1,1\rangle}=2\lambda_{11,11},$ (77)
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&2\gamma_{11,12}\\ 2\gamma_{11,12}&2\gamma_{11,22}\end{bmatrix}$, (78)
$M_{|0,0,0\rangle}=\begin{bmatrix}6\,\lambda_{11,11}&2\sqrt{2}\,\gamma_{11,11}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{11,22}\\ 2\,\gamma_{11,11}&12\,\beta_{11,11}&12\sqrt{2}\,\beta_{11,12}&12\,\beta_{11,22}\\ 2\sqrt{2}\,\gamma_{11,12}&12\sqrt{2}\,\beta_{11,12}&24\,\beta_{11,22}&12\sqrt{2}\,\beta_{12,22}\\ 2\,\gamma_{11,22}&12\,\beta_{11,22}&12\sqrt{2}\,\beta_{12,22}&12\,\beta_{22,22}\end{bmatrix}$. (79)
V.5.2 Unitarity Bounds
Again, certain eigenvalues of the scattering matrices are too complicated to be expressed in closed form. However, others can be written analytically, and, for those, the corresponding unitarity bounds are then applied, as:
$\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{4\gamma_{11,12}^{2}+\left(\gamma_{11,11}-\gamma_{11,22}\right){}^{2}}\right|$ $\leq 8\pi,$ (80)
$|2\lambda_{11,11}|$ $\leq 8\pi,$ (81)
In order to compare our results with previous work, we consider the model studied in Ref. Muhlleitner:2020wwk , where the SM is extended by a neutral complex scalar singlet. In their notation, the quartic part of the scalar potential reads
$V_{4}=\frac{\lambda}{4}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\frac{\delta_{2}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)|S|^{2}+\frac{d_{2}}{2}|S|^{4},$ (83)
where  [^5] Equivalently, one could choose $S=\frac{\chi_{2}+i\chi_{1}}{\sqrt{2}}$. $S=\frac{\chi_{1}+i\chi_{2}}{\sqrt{2}}$. Notice that the case with two real singlets in Eq. (76) is more general than the case with one complex singlet in Eq. (83). But, since the complex scalar singlet can be decomposed into two real scalar singlet fields, $\chi_{1}$ and $\chi_{2}$, we find that our results reproduce those of Muhlleitner:2020wwk , when we set $\gamma_{11,12}=\beta_{11,12}=\beta_{12,22}=0$. By matching the terms in Eq. (83) to those in Eq. (76), we derive the mapping between the two notations that is provided in Table 6.
Term Our Notation Notation in Muhlleitner:2020wwk
$(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\frac{\lambda}{4}$
$\chi_{1}^{4}$ $\beta_{11,11}$ $\frac{d_{2}}{16}$
$\chi_{2}^{4}$ $\beta_{22,22}$ $\frac{d_{2}}{16}$
$\chi_{1}^{2}\chi_{2}^{2}$ $6\beta_{11,22}$ $\frac{d_{2}}{8}$
$\chi_{1}^{3}\chi_{2}$ $4\beta_{11,12}$ 0
$\chi_{1}\chi_{2}^{3}$ $4\beta_{12,22}$ 0
$(\Phi_{1}^{\dagger}\Phi_{1})\chi_{1}^{2}$ $\gamma_{11,11}$ $\frac{\delta_{2}}{4}$
$(\Phi_{1}^{\dagger}\Phi_{1})\chi_{2}^{2}$ $\gamma_{11,22}$ $\frac{\delta_{2}}{4}$
$(\Phi_{1}^{\dagger}\Phi_{1})\chi_{1}\chi_{2}$ $2\gamma_{11,12}$ 0
Table 6: Comparison of coupling notation.
V.6 2 Scalar Doublets and 1 Neutral Scalar Singlet with a $\mathbb{Z}_{2}$ symmetry
We now consider a 2HDM with one additional neutral scalar singlet $\chi_{1}$. We impose a $\mathbb{Z}_{2}$ symmetry that acts on the fields as
$\Phi_{1}\to\Phi_{1},\qquad\Phi_{2}\to-\Phi_{2},\qquad\chi_{1}\to\chi_{1}.$ (84)
Under this symmetry, and using Eqs. (5), (7), and (9), the scalar potential simplifies to:
$V_{4}=\$$\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+2\lambda_{12,21}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)$
$+\left[\lambda_{12,12}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\text{h.c.}\right]+\beta_{11,11}\chi_{1}^{4}+\gamma_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{1}^{2}+\gamma_{22,11}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\chi_{1}^{2}.$ (85)
V.6.1 Scattering Matrices
The set of non-zero scattering matrices for this model reads
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&0\\ 0&2\gamma_{22,11}\end{bmatrix}$, (86)
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda^{*}_{12,12}&0\\ 2\lambda_{12,12}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (87)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda^{*}_{12,12}\\ 0&0&2\lambda_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (88)
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&4\lambda_{11,22}+2\lambda_{12,21}&2\sqrt{2}\,\gamma_{11,11}&0&0\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda_{22,22}&2\sqrt{2}\,\gamma_{22,11}&0&0\\ 2\gamma_{11,11}&2\gamma_{22,11}&12\beta_{11,11}&0&0\\ 0&0&0&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,12}\\ 0&0&0&6\lambda^{*}_{12,12}&2\lambda_{11,22}+4\lambda_{12,21}\end{bmatrix}$, (89)
$M_{|1,1,0\rangle}=2\left(\lambda_{11,22}-\lambda_{12,21}\right).$ (90)
V.6.2 Unitarity Bounds
Some eigenvalues are too complicated to be written in closed form. Therefore, we show only the results of imposing partial-wave unitarity on the remaining eigenvalues of the zero partial-wave amplitude matrix,
$|2\gamma_{11,11}|$ $\leq 8\pi,$ (91)
$|2\gamma_{22,11}|$ $\leq 8\pi,$ (92)
$2|\lambda_{11,22}\pm\lambda_{12,21}|$ $\leq 8\pi,$ (93)
$2\left|\lambda_{11,22}\pm|\lambda_{12,12}|\right|$ $\leq 8\pi,$ (94)
$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4\lambda_{12,21}^{2}}\right|$ $\leq 8\pi,$ (95)
$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq 8\pi.$ (96)
V.7 2 Scalar Doublets with a $\mathbb{Z}_{2}$ symmetry + 2 Neutral Scalar Singlets with a $\mathbb{Z^{\prime}}_{2}$ symmetry
We now consider a model with two $SU(2)$ doublets and two neutral scalar singlets. We impose two independent $\mathbb{Z}_{2}$ symmetries that act on these fields as
$\mathbb{Z}_{2}:$ $\Phi_{1}\to\Phi_{1}\quad\Phi_{2}\to-\Phi_{2},\quad\chi_{1}\to\chi_{1},\quad\chi_{2}\to\chi_{2},$ (97)
$\mathbb{Z}_{2}^{\prime}:$ $\Phi_{1}\to\Phi_{1}\quad\Phi_{2}\to\Phi_{2},\quad\chi_{1}\to-\chi_{1},\quad\chi_{2}\to-\chi_{2},$ (98)
Under these symmetries, the quartic part of the scalar potential becomes:
$V_{4}$ $=\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+2\lambda_{12,21}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)$
$\quad+\left[\lambda_{12,12}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\text{h.c.}\right]+\beta_{11,11}\,\chi_{1}^{4}+\beta_{22,22}\,\chi_{2}^{4}+6\beta_{11,22}\,\chi_{1}^{2}\chi_{2}^{2}+4\beta_{11,12}\,\chi_{1}^{3}\chi_{2}+4\beta_{12,22}\,\chi_{2}^{3}\chi_{1}$
$\quad+\gamma_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{1}^{2}+\gamma_{22,11}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\chi_{1}^{2}+\gamma_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{2}^{2}+2\gamma_{11,12}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\chi_{1}\chi_{2}$
$\quad+2\gamma_{22,12}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\chi_{1}\chi_{2}+\gamma_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\chi_{2}^{2},$ (99)
where we have simplified the quartic couplings using the relations in Eqs. (5) and (7).
V.7.1 Scattering Matrices
The set of non-zero scattering matrices for this model reads
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&2\gamma_{11,12}&0&0\\ 2\gamma_{11,12}&2\gamma_{11,22}&0&0\\ 0&0&2\gamma_{22,11}&2\gamma_{22,12}\\ 0&0&2\gamma_{22,12}&2\gamma_{22,22}\end{bmatrix}$, (100)
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda^{*}_{12,12}&0\\ 2\lambda_{12,12}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (101)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda^{*}_{12,12}\\ 0&0&2\lambda_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (102)
$M_{|1,1,0\rangle}=2\left(\lambda_{11,22}-\lambda_{12,21}\right),$ (103)
$M_{|0,0,0\rangle}=\textrm{blkdiag}(A,B)$ (104)
where, here and henceforth, $\textrm{blkdiag}(A,B,\dots)$ refers to a block diagonal matrix, whose entries are the matrices $A,B,\dots$. The matrices $A$ and $B$ in Eq. (104) are given, respectively, by
$A=\begin{bmatrix}24\,\beta_{11,22}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{22,12}&12\sqrt{2}\,\beta_{11,12}&12\sqrt{2}\,\beta_{12,22}\\ 2\sqrt{2}\,\gamma_{11,12}&6\,\lambda_{11,11}&4\,\lambda_{11,22}+2\,\lambda_{12,21}&2\sqrt{2}\,\gamma_{11,11}&2\sqrt{2}\,\gamma_{11,22}\\ 2\sqrt{2}\,\gamma_{22,12}&4\,\lambda_{11,22}+2\,\lambda_{12,21}&6\,\lambda_{22,22}&2\sqrt{2}\,\gamma_{22,11}&2\sqrt{2}\,\gamma_{22,22}\\ 12\sqrt{2}\,\beta_{11,12}&2\,\gamma_{11,11}&2\,\gamma_{22,11}&12\,\beta_{11,11}&12\,\beta_{11,22}\\ 12\sqrt{2}\,\beta_{12,22}&2\,\gamma_{11,22}&2\,\gamma_{22,22}&12\,\beta_{11,22}&12\,\beta_{22,22}\end{bmatrix}$, (105)
$B=\begin{bmatrix}2\,\lambda_{11,22}+4\,\lambda_{12,21}&6\,\lambda_{12,12}\\ 6\,\lambda_{12,12}^{*}&2\,\lambda_{11,22}+4\,\lambda_{12,21}\end{bmatrix}$. (106)
V.7.2 Unitarity Bounds
As several eigenvalues are too involved to express analytically, we show only the results of imposing partial-wave unitarity on the remaining eigenvalues of the $s$-wave amplitude matrix, that can be computed explicitly,
$2\left|\lambda_{11,22}\pm\lambda_{12,21}\right|$ $\leq$ $8\pi,$ (107)
$2\left|\lambda_{11,22}\pm|\lambda_{12,12}|\right|$ $\leq$ $8\pi,$ (108)
$2\left|\lambda_{11,22}+2\lambda_{12,21}\pm 3|\lambda_{12,12}|\right|$ $\leq$ $8\pi,$ (109)
$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4\lambda_{12,21}^{2}}\right|$ $\leq$ $8\pi,$ (110)
$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq$ $8\pi,$ (111)
$\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{\gamma_{11,11}^{2}+4\gamma_{11,12}^{2}-2\gamma_{11,11}\gamma_{11,22}+\gamma_{11,22}^{2}}\right|$ $\leq$ $8\pi,$ (112)
$\left|\gamma_{22,11}+\gamma_{22,22}\pm\sqrt{\gamma_{22,11}^{2}+4\gamma_{22,12}^{2}-2\gamma_{22,11}\gamma_{22,22}+\gamma_{22,22}^{2}}\right|$ $\leq$ $8\pi.$ (113)
Term Our Notation Notation of Boto:2025mmn
$(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\tfrac{\lambda_{1}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{2})^{2}$ $\lambda_{22,22}$ $\tfrac{\lambda_{2}}{2}$
$(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})$ $2\lambda_{11,22}$ $\lambda_{3}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})$ $2\lambda_{12,21}$ $\lambda_{4}$
$(\Phi_{1}^{\dagger}\Phi_{2})^{2}$ $\lambda_{12,12}$ $\lambda_{5}$
$(\Phi_{2}^{\dagger}\Phi_{1})^{2}$ $\lambda_{12,12}^{*}$ $\lambda_{5}$
$\chi_{1}^{4}$ $\beta_{11,11}$ $\tfrac{\lambda_{6}}{8}$
$\chi_{2}^{4}$ $\beta_{22,22}$ $\tfrac{\lambda_{9}}{8}$
$\chi_{1}^{2}\chi_{2}^{2}$ $6\beta_{11,22}$ $\tfrac{\lambda_{10}}{4}$
$\chi_{1}^{3}\chi_{2}$ $4\beta_{11,12}$ $\tfrac{\lambda_{13}}{6}$
$\chi_{1}\chi_{2}^{3}$ $4\beta_{12,22}$ $\tfrac{\lambda_{14}}{6}$
$(\Phi_{1}^{\dagger}\Phi_{1})\chi_{1}^{2}$ $\gamma_{11,11}$ $\tfrac{\lambda_{7}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{2})\chi_{1}^{2}$ $\gamma_{22,11}$ $\tfrac{\lambda_{8}}{2}$
$(\Phi_{1}^{\dagger}\Phi_{1})\chi_{2}^{2}$ $\gamma_{11,22}$ $\tfrac{\lambda_{11}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{2})\chi_{2}^{2}$ $\gamma_{22,22}$ $\tfrac{\lambda_{12}}{2}$
$(\Phi_{1}^{\dagger}\Phi_{1})\chi_{1}\chi_{2}$ $2\gamma_{11,12}$ $\tfrac{\lambda_{15}}{2}$
$(\Phi_{2}^{\dagger}\Phi_{2})\chi_{1}\chi_{2}$ $2\gamma_{22,12}$ $\tfrac{\lambda_{16}}{2}$
Table 7: Comparison of coupling notation.
We verified our results against those reported in Boto:2025mmn and found them to be in agreement. Table 7 summarizes the correspondence between the couplings used in our work and those of Boto:2025mmn .
V.8 2 Scalar Doublets + 1 Neutral Scalar Singlet + 1 Charged Scalar Singlet
For a model with two $SU(2)$ scalar doublets $\Phi_{1}$ and $\Phi_{2}$, one neutral scalar singlet $\chi_{1}$ and one charged scalar singlet $\varphi^{+}_{1}$, the quartic part of the scalar potential can be written as
$V_{4}=\$$\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+2\lambda_{12,21}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)$
$+\left[\lambda_{12,12}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+2\lambda_{11,12}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{1}^{\dagger}\Phi_{2}\right)+2\lambda_{12,22}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+\text{h.c.}\right]$
$+\alpha_{11,11}\left(\varphi_{1}^{+}\varphi_{1}^{-}\right)^{2}+\beta_{11,11}\left(\chi_{1}^{2}\right)^{2}$
$+\delta_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\varphi_{1}^{+}\varphi_{1}^{-}\right)+\delta_{12,11}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\varphi_{1}^{+}\varphi_{1}^{-}\right)+\delta_{12,11}^{*}\left(\Phi_{2}^{\dagger}\Phi_{1}\right)\left(\varphi_{1}^{+}\varphi_{1}^{-}\right)+\delta_{22,11}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\left(\varphi_{1}^{+}\varphi_{1}^{-}\right)$
$+\gamma_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\chi_{1}^{2}\right)+\gamma_{12,11}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\chi_{1}^{2}\right)+\gamma_{12,11}^{*}\left(\Phi_{2}^{\dagger}\Phi_{1}\right)\left(\chi_{1}^{2}\right)+\gamma_{22,11}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\left(\chi_{1}^{2}\right)+\zeta_{11,11}\left(\varphi_{1}^{+}\varphi_{1}^{-}\right)\left(\chi_{1}^{2}\right)$
$+\left[\kappa_{12,11}\left(\Phi_{1}^{T}\sigma_{2}\Phi_{2}\right)\left(\varphi_{1}^{-}\chi_{1}\right)-\kappa_{12,11}\left(\Phi_{2}^{T}\sigma_{2}\Phi_{1}\right)\left(\varphi_{1}^{-}\chi_{1}\right)+\mathrm{h.c.}\right],$ (114)
where we have used the relations in Eqs. (5), (8), (9), and (11) to get rid of redundant terms. This is the simplest model featuring the $\kappa_{ab,cd}$ couplings from Eq. (4).
V.8.1 Scattering Matrices
The set of non-zero scattering matrices are:
$M_{|2,2,0\rangle}=2\alpha_{11,11},$ (115)
$M_{|2,\frac{3}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}^{*}\\ \delta_{12,11}&\delta_{22,11}\end{bmatrix}$, (116)
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}&0&-2i\kappa^{*}_{12,11}\\ \delta^{*}_{12,11}&\delta_{22,11}&2i\kappa^{*}_{12,11}&0\\ 0&-2i\kappa_{12,11}&2\gamma_{11,11}&2\gamma^{*}_{12,11}\\ 2i\kappa_{12,11}&0&2\gamma_{12,11}&2\gamma_{22,11}\end{bmatrix}$, (117)
$M_{|1,1,0\rangle}=\begin{bmatrix}2\lambda_{11,22}-2\lambda_{12,21}&2i\sqrt{2}\kappa^{*}_{12,11}\\ -2i\sqrt{2}\kappa^{*}_{12,11}&2\zeta_{11,11}\end{bmatrix}$, (118)
$M_{|1,1,2\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{11,12}^{*}&2\lambda_{12,12}^{*}\\ 2\sqrt{2}\lambda_{11,12}&2\lambda_{11,22}+2\lambda_{12,21}&2\sqrt{2}\lambda_{12,22}^{*}\\ 2\lambda_{12,12}&2\sqrt{2}\lambda_{12,22}&2\lambda_{22,22}\end{bmatrix}$, (119)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{11,12}&2\lambda_{11,12}^{*}&2\lambda_{12,21}\\ 2\lambda_{11,12}^{*}&2\lambda_{11,22}&2\lambda_{12,12}^{*}&2\lambda_{12,22}^{*}\\ 2\lambda_{11,12}&2\lambda_{12,12}&2\lambda_{11,22}&2\lambda_{12,22}\\ 2\lambda_{12,21}&2\lambda_{12,22}&2\lambda_{12,22}^{*}&2\lambda_{22,22}\end{bmatrix}$, (120)
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&6\lambda^{*}_{11,12}&6\lambda_{11,12}&4\lambda_{11,22}+2\lambda_{12,21}&\sqrt{2}\delta_{11,11}&2\sqrt{2}\gamma_{11,11}\\ 6\lambda_{11,12}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,12}&6\lambda_{12,22}&\sqrt{2}\delta_{12,11}&2\sqrt{2}\gamma_{12,11}\\ 6\lambda^{*}_{11,12}&6\lambda^{*}_{12,12}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda^{*}_{12,22}&\sqrt{2}\delta^{*}_{12,11}&2\sqrt{2}\gamma^{*}_{12,11}\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda^{*}_{12,22}&6\lambda_{12,22}&6\lambda_{22,22}&\sqrt{2}\delta_{22,11}&2\sqrt{2}\gamma_{22,11}\\ \sqrt{2}\delta_{11,11}&\sqrt{2}\delta^{*}_{12,11}&\sqrt{2}\delta_{12,11}&\sqrt{2}\delta_{22,11}&4\alpha_{11,11}&\sqrt{2}\zeta_{11,11}\\ 2\gamma_{11,11}&2\gamma^{*}_{12,11}&2\gamma_{12,11}&2\gamma_{22,11}&\sqrt{2}\zeta_{11,11}&12\beta_{11,11}\end{bmatrix}$. (121)
V.8.2 Unitarity Bounds
Since certain eigenvalues are too complex to evaluate analytically, unitarity constraints are shown only for the remaining eigenvalues of the zero partial-wave amplitude matrix. These read
$\left|\zeta_{11,11}+\lambda_{11,22}-\lambda_{12,21}\pm\sqrt{\left(-\zeta_{11,11}-\lambda_{11,22}+\lambda_{12,21}\right)^{2}-4\left(\zeta_{11,11}\lambda_{11,22}-\zeta_{11,11}\lambda_{12,21}-2|\kappa_{12,11}|^{2}\right)}\right|$ $\leq 8\pi,$ (122)
$\frac{1}{2}\left|\delta_{11,11}+\delta_{22,11}\pm\sqrt{\left(\delta_{11,11}-\delta_{22,11}\right)^{2}+4|\delta_{12,11}|^{2}}\right|$ $\leq 8\pi,$ (123)
$\left|2~\alpha_{11,11}\right|$ $\leq 8\pi.$ (124)
V.9 $\mathbb{Z}_{3}$-Symmetric 3HDM
Now, we consider a model with three $SU(2)$ doublets $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$. Under a $\mathbb{Z}_{3}$ symmetry, the three doublets transform as
$\Phi_{1}\to\Phi_{1},\qquad\Phi_{2}\to e^{i\frac{2\pi}{3}}\Phi_{2},\qquad\Phi_{3}\to e^{i\frac{4\pi}{3}}\Phi_{3}.$ (125)
Under this symmetry, the potential becomes
$V_{4}=$ $\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{33,33}\left(\Phi_{3}^{\dagger}\Phi_{3}\right)^{2}$ (126)
$+2\lambda_{11,22}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+2\lambda_{11,33}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{3}^{\dagger}\Phi_{3}\right)+2\lambda_{22,33}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)\left(\Phi_{3}^{\dagger}\Phi_{3}\right)$
$+2\lambda_{12,21}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{1}\right)+2\lambda_{13,31}\left(\Phi_{1}^{\dagger}\Phi_{3}\right)\left(\Phi_{3}^{\dagger}\Phi_{1}\right)+2\lambda_{23,32}\left(\Phi_{2}^{\dagger}\Phi_{3}\right)\left(\Phi_{3}^{\dagger}\Phi_{2}\right)$
$+2\left[\lambda_{12,13}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{1}^{\dagger}\Phi_{3}\right)+\lambda_{13,23}\left(\Phi_{1}^{\dagger}\Phi_{3}\right)\left(\Phi_{2}^{\dagger}\Phi_{3}\right)+\lambda_{12,32}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{3}^{\dagger}\Phi_{2}\right)+\textrm{h.c.}\right],$
where $\Phi_{1}$$\Phi_{2}$ and $\Phi_{3}$ denote the three scalar doublets.
V.9.1 Scattering Matrices
The scattering matrices are all block diagonal. We find
$M_{|2,1,1\rangle}=\textrm{blkdiag}(A,B,C),$ (127)
where
$A$ $=$ $\begin{bmatrix}2\lambda_{11,33}+2\lambda_{13,31}&2\sqrt{2}\lambda^{*}_{12,32}\\[2.84526pt] 2\sqrt{2}\lambda_{12,32}&2\lambda_{22,22}\end{bmatrix}$, (128)
$B$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda^{*}_{12,13}\\[2.84526pt] 2\sqrt{2}\lambda_{12,13}&2\lambda_{22,33}+2\lambda_{23,32}\end{bmatrix}$, (129)
$C$ $=$ $\begin{bmatrix}2\lambda_{33,33}&2\sqrt{2}\lambda_{13,23}\\[2.84526pt] 2\sqrt{2}\lambda^{*}_{13,23}&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (130)
$M_{|1,0,1\rangle}=\textrm{blkdiag}(D,E,F),$ (131)
where
$D$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{13,23}&2\lambda^{*}_{12,32}\\ 2\lambda^{*}_{13,23}&2\lambda_{11,33}&2\lambda^{*}_{12,13}\\ 2\lambda_{12,32}&2\lambda_{12,13}&2\lambda_{11,22}\end{bmatrix}$, (132)
$E$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&2\lambda_{13,31}\\ 2\lambda_{12,21}&2\lambda_{22,22}&2\lambda_{23,32}\\ 2\lambda_{13,31}&2\lambda_{23,32}&2\lambda_{33,33}\end{bmatrix}$, (133)
$F$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{12,32}&2\lambda^{*}_{13,23}\\ 2\lambda^{*}_{12,32}&2\lambda_{11,22}&2\lambda^{*}_{12,13}\\ 2\lambda_{13,23}&2\lambda_{12,13}&2\lambda_{11,33}\end{bmatrix}$, (134)
$M_{|0,0,0\rangle}=\textrm{blkdiag}(G,H,I),$ (135)
where
$G$ $=$ $\begin{bmatrix}2\lambda_{22,33}+4\lambda_{23,32}&6\lambda^{*}_{13,23}&6\lambda_{12,32}\\ 6\lambda_{13,23}&2\lambda_{11,33}+4\lambda_{13,31}&6\lambda_{12,13}\\ 6\lambda^{*}_{12,32}&6\lambda^{*}_{12,13}&2\lambda_{11,22}+4\lambda_{12,21}\end{bmatrix}$, (136)
$H$ $=$ $\begin{bmatrix}6\lambda_{11,11}&4\lambda_{11,22}+2\lambda_{12,21}&4\lambda_{11,33}+2\lambda_{13,31}\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda_{22,22}&4\lambda_{22,33}+2\lambda_{23,32}\\ 4\lambda_{11,33}+2\lambda_{13,31}&4\lambda_{22,33}+2\lambda_{23,32}&6\lambda_{33,33}\end{bmatrix}$, (137)
$I$ $=$ $\begin{bmatrix}2\lambda_{22,33}+4\lambda_{23,32}&6\lambda^{*}_{12,32}&6\lambda_{13,23}\\ 6\lambda_{12,32}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,13}\\ 6\lambda^{*}_{13,23}&6\lambda^{*}_{12,13}&2\lambda_{11,33}+4\lambda_{13,31}\end{bmatrix}$, (138)
and
$M_{|1,1,0\rangle}=\begin{bmatrix}2(\lambda_{11,22}-\lambda_{12,21})&0&0\\ 0&2(\lambda_{11,33}-\lambda_{13,31})&0\\ 0&0&2(\lambda_{22,33}-\lambda_{23,32})\end{bmatrix}$. (139)
V.9.2 Unitarity Bounds
We show the results of partial wave unitarity bounds to the eigenvalues that can be computed analytically,
$2|\lambda_{11,22}-\lambda_{12,21}|$ $\leq 8\pi,$ (140)
$2|\lambda_{11,33}-\lambda_{13,31}|$ $\leq 8\pi,$ (141)
$2|\lambda_{22,33}-\lambda_{23,32}|$ $\leq 8\pi,$ (142)
$\left|\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\pm\sqrt{\left(\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\right)^{2}-4\left(\lambda_{11,11}\lambda_{22,33}+\lambda_{11,11}\lambda_{23,32}-2|\lambda_{12,13}|^{2}\right)}\right|$ $\leq 8\pi,$ (143)
$\left|\lambda_{11,33}+\lambda_{13,31}+\lambda_{22,22}\pm\sqrt{\left(\lambda_{11,33}+\lambda_{13,31}+\lambda_{22,22}\right)^{2}-4\left(\lambda_{11,33}\lambda_{22,22}+\lambda_{13,31}\lambda_{22,22}-2|\lambda_{12,32}|^{2}\right)}\right|$ $\leq 8\pi,$ (144)
$\left|\lambda_{11,22}+\lambda_{12,21}+\lambda_{33,33}\pm\sqrt{\left(\lambda_{11,22}+\lambda_{12,21}+\lambda_{33,33}\right)^{2}-4\left(\lambda_{11,22}\lambda_{33,33}+\lambda_{12,21}\lambda_{33,33}-2\left|\lambda_{13,23}\right|^{2}\right)}\right|$ $\leq 8\pi.$ (145)
Term Our Notation Notation of Bento:2017eti
$(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $r_{1}$
$(\Phi_{2}^{\dagger}\Phi_{2})^{2}$ $\lambda_{22,22}$ $r_{2}$
$(\Phi_{3}^{\dagger}\Phi_{3})^{2}$ $\lambda_{33,33}$ $r_{3}$
$(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})$ $2\lambda_{11,22}$ $2r_{4}$
$(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{3}^{\dagger}\Phi_{3})$ $2\lambda_{11,33}$ $2r_{5}$
$(\Phi_{2}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{3})$ $2\lambda_{22,33}$ $2r_{6}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})$ $2\lambda_{12,21}$ $2r_{7}$
$(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{1})$ $2\lambda_{13,31}$ $2r_{8}$
$(\Phi_{2}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{2})$ $2\lambda_{23,32}$ $2r_{9}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{1}^{\dagger}\Phi_{3})$ $2\lambda_{12,13}$ $2c_{4}$
$(\Phi_{2}^{\dagger}\Phi_{1})(\Phi_{3}^{\dagger}\Phi_{1})$ $2\lambda_{12,13}^{\ast}$ $2c_{4}^{\ast}$
$(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{2}^{\dagger}\Phi_{3})$ $2\lambda_{13,23}$ $2c_{11}$
$(\Phi_{3}^{\dagger}\Phi_{1})(\Phi_{3}^{\dagger}\Phi_{2})$ $2\lambda_{13,23}^{\ast}$ $2c_{11}^{\ast}$
$(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{2})$ $2\lambda_{12,32}$ $2c_{12}$
$(\Phi_{2}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{3})$ $2\lambda_{12,32}^{\ast}$ $2c_{12}^{\ast}$
Table 8: Comparison of coupling notation.
We compared our results with those presented in Bento:2017eti , and found them to be consistent. Table 8 provides a summary of the correspondence between the quartic couplings used in our study and those in Bento:2017eti .
VI Overview
Almost all models addressing the outstanding issues in the SM include extra $SU(2)$ singlet and/or doublet scalars. In particular, many models addressing the dark matter problem include extra neutral singlet scalars. Such models must be subject to theoretical constraints, even before a simulation starts. Those constraints include boundedness from below, nonexistence of lower lying alternative vaccua, and the unitarity bounds on $2\rightarrow 2$ scattering. In this paper we address fully the perturbative partial-wave unitarity for the tree-level scattering matrix in models with any number of scalar doublets, neutral singlets, and/or charged singlets. Enforcing the correct high-energy behavior, provides bounds on the quartic couplings, freeing us from defining the exact nature of the quadratic and cubic couplings.
In contrast, if one one wishes to turn such bounds on restrictions on masses, mixing angles and other directly observable quantities, then one must define the full theory. That is, one must define the quadratic and cubic couplings, a specific vacuum, and the mass matrices must be dully diagonalized. After this, one would strife to invert the relations, turning the unitarity bounds on the quartic couplings into restrictions on combinations of masses and mixing angles. This is possible in simple models, but increasingly more difficult as the number of fields increases. In contrast, although the matrices get larger, our limits on quartic couplings are always applicable, at least numerically.
We classify the states by the conserved quantum numbers $Q$, $Y$, and $T$ and show that, once one restricts oneself to the minimal set of states providing all inequivalent bounds, the quantum number $Q$ is redundant. We also discuss examples where the existence of extra symmetries allows for the inclusion of further quantum numbers in the basis, thus greatly simplifying the matrices.
We introduce the Mathematica notebook BounDS that automatically calculates the quartic part of the potential and the scattering matrices and their eigenvalues, for any model with any symmetries (discrete or continuous, abelian or non-abelian). We present results for a variety of particular models, and compare with the literature, when available. Our aim is to help provide complete simulations of models beyond the SM with a necessary and very powerful tool in parameter restriction.
Acknowledgments
We are grateful to Rui Santos for reading and commenting on this document. This work is supported in part by the Portuguese Fundação para a Ciência e Tecnologia (FCT) through the PRR (Recovery and Resilience Plan), within the scope of the investment ”RE-C06-i06 - Science Plus Capacity Building”, measure ”RE-C06-i06.m02 - Reinforcement of financing for International Partnerships in Science, Technology and Innovation of the PRR”, under the project with reference 2024.01362.CERN. The work of the authors is also supported by FCT under Contracts UIDB/00777/2020, and UIDP/00777/2020. The FCT projects are partially funded through POCTI (FEDER), COMPETE, QREN, and the EU. A.M. was additionally supported by FCT with PhD Grant No. 2024.01340.BD.
Appendix A Redundant scattering matrices
In any scattering process, conserved quantum numbers constrain the possible initial and final states. In a globally symmetric $SU(2)\times U(1)$ quantum field theory, initial states with definite electric charge $Q$ and hypercharge $Y$ can only scatter into states with the same $Q$ and $Y$. That is because these quantum numbers are protected by the $SU(2)\times U(1)$ symmetry and are, thus, conserved. For this reason, in building $2\to 2$ scattering matrices, Ref. Bento:2017eti labeled two-particle states by $|Q,Y\rangle$. For models like the one defined in Section II, we list in Table. 9 all possible two-particle states.
$|Q,Y\rangle$ State Conditions Dimensionality
$|2,2\rangle$ $\varphi^{+}_{i}\varphi^{+}_{j}$ $i\leq j$ $n_{c}(n_{c}+1)/2$
$|2,\frac{3}{2}\rangle$ $\phi^{+}_{i}\varphi^{+}_{j}$ $n_{D}n_{c}$
$|2,1\rangle$ $\phi^{+}_{i}\phi^{+}_{j}$ $i\leq j$ $n_{D}(n_{D}+1)/2$
$|1,\frac{3}{2}\rangle$ $\phi^{0}_{i}\varphi^{+}_{j}$ $n_{D}n_{c}$
$|1,1\rangle$ $\left\{\phi^{+}_{i}\,\phi^{0}_{j},\ \varphi^{+}_{i}\chi_{j}\right\}$ $n_{D}^{2}+n_{n}n_{c}$
$|1,\frac{1}{2}\rangle$ $\left\{\phi^{+}_{i}\chi_{j},\ \phi^{0*}_{i}\varphi^{+}_{j}\right\}$ $n_{D}(n_{n}+n_{c})$
$|1,0\rangle$ $\phi^{+}_{i}\phi^{0*}_{j}$ $n_{D}^{2}$
$|0,1\rangle$ $\phi^{0}_{i}\phi^{0}_{j}$ $i\leq j$ $n_{D}(n_{D}+1)/2$
$|0,\frac{1}{2}\rangle$ $\left\{\phi^{0}_{i}\chi_{j},\ \phi^{-}_{i}\varphi^{+}_{j}\right\}$ $n_{D}(n_{n}+n_{c})$
$|0,0\rangle$ $\left\{\phi_{i}^{+}\phi_{j}^{-},\phi_{i}^{0}\phi_{j}^{0*},\ \varphi_{i}^{+}\varphi_{j}^{-},\ \chi_{i}\chi_{j}\right\}$ $\{\,\textrm{---}\,,\,\textrm{---}\,,\,\textrm{---}\,,\,i\leq j\}$ $2n_{D}^{2}+n_{c}^{2}+n_{n}(n_{n}+1)/2$
Table 9: Basis of two-particle states labeled by $|Q,Y\rangle$. This table includes all states, some of which provide redundant information.
However, not all scattering amplitudes are independent. Notice that
$\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\delta_{ca,db},$ (146)
$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\phi^{0}_{b}\to\phi^{0}_{c}\phi^{0}_{d}\right]=2\lambda_{ca,db}+2\lambda_{da,cb},$ (147)
$\mathcal{M}\left[\phi^{+}_{a}\chi_{b}\to\phi^{+}_{c}\chi_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\chi_{b}\to\phi^{0}_{c}\chi_{d}\right]=2\gamma_{ca,bd},$ (148)
$\mathcal{M}\left[\phi^{+}_{a}\chi_{b}\to\phi^{0*}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\chi_{b}\to\phi^{-}_{c}\varphi^{+}_{d}\right]=2i\kappa_{ca,db},$ (149)
$\mathcal{M}\left[\phi^{0*}_{a}\varphi^{+}_{b}\to\phi^{0*}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{-}_{a}\varphi^{+}_{b}\to\phi^{-}_{c}\varphi^{+}_{d}\right]=\delta_{ac,db}.$ (150)
Therefore, perturbative unitarity bounds obtained from scatterings involving $|1,\frac{3}{2}\rangle$, $|0,1\rangle$ , and $|0,\frac{1}{2}\rangle$ are redundant because they are identical to those derived from $|2,\frac{3}{2}\rangle$, $|2,1\rangle$ , and $|1,\frac{1}{2}\rangle$ , respectivelly.
In any scattering involving $SU(2)$ doublets, total isospin $T$ must also be conserved, For this reason, two-particle states should be further labeled by $|Q,Y,T\rangle$, with $T=0,1$ in this model. Using Clebsch-Gordan coefficients, we can split the two-particle state $\phi^{+}_{i}\phi^{0}_{j}$ in $|1,1\rangle$ into
$|1,1,0\rangle:$ $\phi^{+}_{[i}\,\phi^{0}_{j]}\,\equiv\,\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{0}_{j}-\phi^{+}_{j}\,\phi^{0}_{i}\right),$ (151)
$|1,1,1\rangle:$ $\phi^{+}_{(i}\,\phi^{0}_{j)}\,\equiv\,\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{0}_{j}+\phi^{+}_{j}\,\phi^{0}_{i}\right).$ (152)
In the same manner, we can split the two-particle state $\phi^{+}_{i}\phi^{-}_{j}$ in $|0,0\rangle$ into
$|0,0,0\rangle:$ $\Phi_{i}\,\Phi_{j}^{*}\,\equiv\,\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{-}_{j}+\phi^{0}_{i}\,\phi^{0*}_{j}\right),$ (153)
$|0,0,1\rangle:$ $\overline{\Phi_{i}\,\Phi_{j}^{*}}\,\equiv\,\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{-}_{j}-\phi^{0}_{i}\,\phi^{0*}_{j}\right).$ (154)
These redefinitions of states lead, once again, to redundant scattering matrices because
$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{(a}\,\phi^{0}_{b)}\to\phi^{+}_{(c}\,\phi^{0}_{d)}\right]=2\lambda_{ca,db}+2\lambda_{da,cb},$ (155)
$\mathcal{M}\left[\phi^{+}_{a}\phi^{0*}_{b}\to\phi^{+}_{c}\phi^{0*}_{d}\right]=\mathcal{M}\left[\overline{\Phi_{a}\Phi_{b}^{*}}\to\overline{\Phi_{c}\Phi_{d}^{*}}\right]=2\lambda_{ca,bd}.$ (156)
We therefore conclude that it is sufficient to apply the partial wave unitarity bounds to the scattering matrices built out of the states listed in Table 1, i.e. the states
$\left|Q,Y,T\right\rangle\quad=\quad\left|2,2,0\right\rangle,\ \left|2,\frac{3}{2},\frac{1}{2}\right\rangle,\ \left|2,1,1\right\rangle,\ \left|1,1,0\right\rangle,\ \left|1,\frac{1}{2},\frac{1}{2}\right\rangle,\ \left|1,0,1\right\rangle,\ \left|0,0,0\right\rangle.$ (157)
Appendix B The Standard Model example
Labeling states by $Q$ and $Y$
If we classify the states solely by their electric charge $Q$ and hypercharge $Y$, the SM requires computing the scattering matrices for the states listed in Table 10.
$|Q,Y\rangle$ State Dimensionality
$|2,1\rangle$ $\phi^{+}_{1}\phi^{+}_{1}$ 1
$|1,1\rangle$ $\phi^{+}_{1}\phi^{0}_{1}$ 1
$|1,0\rangle$ $\phi^{+}_{1}\phi^{0*}_{1}$ 1
$|0,1\rangle$ $\phi^{0}_{1}\phi^{0}_{1}$ 1
$|0,0\rangle$ $\left\{\phi^{+}_{1}\phi^{-}_{1},\ \phi^{0}_{1}\phi^{0*}_{1}\right\}$ 2
Table 10: Basis of two-particle states labelled by $|Q,Y\rangle$.
The corresponding scattering matrices are
$M_{|0,1\rangle}=M_{|1,0\rangle}=M_{|1,1\rangle}=M_{|2,1\rangle}$ $=2\lambda_{11,11},$ (158)
$M_{|0,0\rangle}$ $=\begin{bmatrix}4\lambda_{11,11}&2\lambda_{11,11}\\ 2\lambda_{11,11}&4\lambda_{11,11}\end{bmatrix}$. (159)
Including Total Isospin
We now refine the classification by also incorporating the total isospin $T$. The two-particulate states with definite $Q$, $Y$, and $T$ are, thus, the ones listed in Table 11.
$|Q,Y,T\rangle$ State Dimensionality
$|2,1,1\rangle$ $\phi^{+}_{1}\phi^{+}_{1}$ 1
$|1,1,1\rangle$ $\phi^{+}_{(1}\phi^{0}_{1)}$ 1
$|1,0,1\rangle$ $\phi^{+}_{1}\phi^{0*}_{1}$ 1
$|0,0,1\rangle$ $\overline{\Phi_{1}\Phi_{1}^{*}}$ 1
$|0,0,0\rangle$ $\Phi_{1}\Phi_{1}^{*}$ 1
$|0,1,1\rangle$ $\phi^{0}_{1}\phi^{0}_{1}$ 1
Table 11: Basis of two-particle states labelled by $|Q,Y,T\rangle$.
Let transformation from the old $(Q,Y)$ basis to the new $(Q,Y,T)$ basis be given by
$\begin{bmatrix}\Phi_{1}\Phi_{1}^{*}\\ \overline{\Phi_{1}\Phi_{1}^{*}}\end{bmatrix}=U\,\begin{bmatrix}\phi_{1}^{+}\phi_{1}^{-}\\ \phi_{1}^{0}\phi_{1}^{0*}\end{bmatrix}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\begin{bmatrix}\phi_{1}^{+}\phi_{1}^{-}\\ \phi_{1}^{0}\phi_{1}^{0*}\end{bmatrix}$. (160)
As a result, the scattering matrix for the states $|0,0\rangle$, gets diagonalized as
$(M_{|0,0\rangle})_{\text{new}}=U(M_{|0,0\rangle})_{\text{old}}U^{\dagger}=\begin{bmatrix}6\lambda_{11,11}&0\\ 0&2\lambda_{11,11}\end{bmatrix}$. (161)
Similarly, for the $|1,1\rangle$ states, only the symmetric combination survives
$\phi^{+}_{(1}\phi^{0}_{1)}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\end{bmatrix}\begin{bmatrix}\phi_{1}^{+}\phi_{1}^{0}\\ \phi_{1}^{0}\phi_{1}^{+}\end{bmatrix}$, (162)
meaning
$(M_{|1,1\rangle})_{\text{new}}=2\lambda_{11,11}.$ (163)
Collecting all cases, the scattering eigenvalues are
$M_{|0,0,0\rangle}$ $=6\lambda_{11,11},$ (164)
$M_{|0,0,1\rangle}$ $=M_{|1,1,0\rangle}=M_{|2,1\rangle}=M_{|1,0\rangle}=M_{|0,1\rangle}=2\lambda_{11,11}.$ (165)
The eigenvalues, and therefore the unitarity bounds, follow directly. This demonstrates the advantage of organizing states in the $|Q,Y,T\rangle$ basis: the scattering matrices simplify and, in this case, diagonalize naturally.
Appendix C BounDS
We have developed BounDS, a Mathematica notebook, that automates the process of writing the quartic potential and deriving partial wave unitarity bounds for models within the class of models defined in Section II. Notice that ScannerS Coimbra:2013qq ; Muhlleitner:2020wwk has a tool to calculate the scattering matrices, but the user must introduce the correct potential. Here, the user just introduces the number of fields and their symmetries; our program calculates the correct quartic potential and then the relevant scattering matrices. As stated, the notebook can be downloaded from
https://github.com/andremilagre/BounDS.git
The notebook BounDS is divided into three parts. To illustrate how it works, we consider the $U(1)$-symmetric 2HDM. In step 1, the user specifies the number of fields to include.
nD=2;
nc=0;
nn=0;
If needed, the user can also impose additional symmetries on the fields by specifying the number of symmetries nSym, and the way the symmetries act on the fields. The latter is done by populating the vector Sym[]:
nSym = 1;
Sym[1] = {
$\Phi$[1] -> $\Phi$[1],
$\Phi$[2] -> Exp[I a] $\Phi$[2]
Assume = {a $\in$ Reals};
The list Assume contains information about group theoretical parameters, such as the space they are defined in and relationships between them. For these parameters, the user must use Latin characters. Furthermore, the function Conjugate[] may be used for symmetries that require conjugation of a field. BounDS has been successfully tested with a wide range of symmetry groups, including both discrete and continuous, abelian and non-abelian cases.
After evaluating the cells in step 1, the user can simply run all the cells in step 2. Under the hood, this module does the following operations:
  • Lists the minimal set of linearly independent quartic couplings allowed by the symmetries.
  • Assembles the 7 independent scattering matrices defined in Section III.2.
  • Block-diagonalizes the scattering matrices by swapping rows and columns.
Finally, step 3 is dedicated to visualizing and analyzing the output. By calling Potential4, the user can output the quartic part of the most general scalar potential allowed by the symmetries of the model. Anywhere in the code, the Mathematica function TeXForm can be called to get the source code:
Potential4 //TeXForm
$V_{4}$ $=$ $\lambda_{11,11}\left(\Phi_{1}^{\dagger}\Phi_{1}\right){}^{2}+2\lambda_{11,22}\Phi_{2}^{\dagger}\Phi_{2}\,\Phi_{1}^{\dagger}\Phi_{1}+2\lambda_{12,21}\Phi_{1}^{\dagger}\Phi_{2}\,\Phi_{2}^{\dagger}\Phi_{1}+\lambda_{22,22}\left(\Phi_{2}^{\dagger}\Phi_{2}\right){}^{2}$. (166)
Furthermore, the user may access the basis vector and the corresponding scattering matrix by specifying the value of $Q$, $Y$, and $T$, and evaluating the functions Basis[Q, Y, T] and ScatteringMatrix[Q, Y, T], respectively:
Q = 0;
Y = 0;
T = 0;
Basis[Q, Y, T] //TeXForm
ScatteringMatrix[Q, Y, T] //TeXForm
$\left(\begin{array}[]{cccc}\Phi_{2}\Phi_{2}^{\dagger},&\Phi_{1}\Phi_{1}^{\dagger},&\Phi_{2}\Phi_{1}^{\dagger},&\Phi_{1}\Phi_{2}^{\dagger}\\ \end{array}\right)\,$, (168)
$\begin{bmatrix}6\lambda_{22,22}&2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&0&0\\ 2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&6\lambda_{11,11}&0&0\\ 0&0&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&0\\ 0&0&0&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)\\ \end{bmatrix}$. (169)
Note that the ordering of the two-particle states in the basis vector may differ from the expected convention. This is an artifact of the block-diagonalization routine. However, the corresponding eigenvalues are unaffected by this reordering.
Finally, the list of eigenvalues for a given scattering matrix can be obtained by calling EigenList[Q, Y, T], again, with definite values for $Q$, $Y$, and $T$:
Q = 0;
Y = 0;
T = 0;
EigenList[Q, Y, T] //FullSimplify //TeXForm
[…]
Appendix D $\mathbb{Z}_{2}$-Symmetric 2HDM
In a process involving complex scalars that are singlets under the SM gauge group, there are CP-even and CP-odd components that scatter independently. Furthermore, if there are additional flavour symmetries, each two-particle state must also be labelled by its corresponding charge $S_{\alpha}$, following the idea of Ginzburg:2005dt . Therefore, we can label all states by,
$|Q,Y,T,\mathbf{\text{{CP}},S_{1},S_{2},\cdots}\rangle.$ (175)
Let us consider the $\mathbb{Z}_{2}$-symmetric 2HDM with the two-particle states labeled by $|Q,Y,T,\mathbb{Z}_{2}\rangle$. The basis of states is explicitly given in Table 12, where, in the last two lines, we have used the definition in Eq. (27).
$|Q,Y,T,\mathbb{Z}_{2}\rangle$ State Dimensionality
$|2,1,1,+1\rangle$ $\{\phi^{+}_{1}\phi^{+}_{1},\,\phi^{+}_{2}\phi^{+}_{2}\}$ 2
$|2,1,1,-1\rangle$ $\phi^{+}_{1}\phi^{+}_{2}$ 1
$|1,1,0,-1\rangle$ $\phi^{+}_{[1}\phi^{0}_{2]}$ 1
$|1,0,1,+1\rangle$ $\{\phi^{+}_{1}\phi^{0*}_{1},\,\phi^{+}_{2}\phi^{0*}_{2}\}$ 2
$|1,0,1,-1\rangle$ $\{\phi^{+}_{1}\phi^{0*}_{2},\,\phi^{+}_{2}\phi^{0*}_{1}\}$ 2
$|0,0,0,+1\rangle$ $\{\Phi_{1}\Phi_{1}^{*},\,\Phi_{2}\Phi_{2}^{*}\}$ 2
$|0,0,0,-1\rangle$ $\{\Phi_{1}\Phi_{2}^{*},\,\Phi_{2}\Phi_{1}^{*}\}$ 2
Table 12: Basis of two-particle states labelled by $|Q,Y,T,\mathbb{Z}_{2}\rangle$.
The scattering matrices are thus
[…]
$M_{\left|1,\,1,\,0,\,-1\right\rangle}$ $=$ $2\left(\lambda_{11,22}-\lambda_{12,21}\right),$ (178)
[…]
$M_{\left|0,\,0,\,0,\,+1\right\rangle}$ $=$ $\begin{bmatrix}6\lambda_{11,11}&2\left(2\lambda_{11,22}+\lambda_{12,21}\right)\\ 2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&6\lambda_{22,22}\end{bmatrix}$, (181)
$M_{\left|0,\,0,\,0,\,-1\right\rangle}$ $=$ $\begin{bmatrix}2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,12}\\ 6\lambda_{12,12}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)\end{bmatrix}$. (182)
Using the substitutions in Table 4 we recover the well-known unitarity bounds for the $\mathbb{Z}_{2}$-symmetric 2HDM, shown in Eqs. (63)–(68).
We find that, by labelling states with flavour symmetries in addition to charge, hypercharge, and total isospin, the scattering matrices simplify further and reduce in dimensionality compared to those in Section V.3.1. This reduction greatly facilitates the subsequent calculations, as shown in this appendix.

### v3

@@
-October 17, 2025
+Perturbative unitarity for models with singlet and doublet scalars
 Abstract
-We provide a complete description of unitarity bounds on the gauge-scalar sectors of models with extra $SU(2)$ doublet, neutral singlet, and charged singlet scalars.
+We provide a complete description of perturbative unitarity bounds on the gauge-scalar sectors of models with extra $SU(2)$ doublet, neutral singlet, and charged singlet scalars.
 Such additions are very frequent in models beyond the Standard Model, and, in particular, they are almost universal in models explaining the dark matter problem.
@@
 I Introduction
-The Standard Model (SM) of particle physics has been extremely successful, culminating in the discovery of a scalar particle consistent with the Higgs boson ATLAS:2012yve ; CMS:2012qbp .
-Nevertheless, it is not a complete theory, and extensions beyond the Standard Model (BSM) are required.
-In particular, the SM provides no explanation for dark matter (DM) ParticleDataGroup:2016lqr , whose presence is inferred across a wide range of astronomical scales, from kiloparsecs to the size of the observable Universe.
-Key probes include the rotation curves of spiral galaxies, the study of galaxy clusters, and measurements of the Cosmic Microwave Background Planck:2018vyg ; Cirelli:2024ssz .
+The Standard Model (SM) of particle physics has been extremely successful, culminating in the discovery of a scalar particle consistent with the Higgs boson Aad and others (2012); Chatrchyan and others (2012).
+Nevertheless, it is unlikely to be the final theory, and extensions beyond the Standard Model are widely considered.
+In particular, the SM provides no explanation for dark matter Navas and others (2024), whose presence is inferred across a wide range of astronomical scales, from kiloparsecs to the size of the observable Universe.
+Key probes include the rotation curves of spiral galaxies, the study of galaxy clusters, and measurements of the Cosmic Microwave Background Aghanim and others (2020); Cirelli et al.
+(2024).
 One of the most widely studied scenarios envisions dark matter as consisting of new fundamental particles.
-The simplest scalar extension introduces a real, neutral $SU(2)$ singlet with a global $\mathbb{Z}_{2}$ symmetry to ensure stability Burgess:2000yq ; GAMBIT:2017gge .
-Alternatively, a minimal fermionic option is a Weyl fermion, commonly referred to as a sterile neutrino Cirelli:2024ssz .
-Other proposed models include the Inert Doublet Model, which extends the Higgs sector with a dark doublet Ilnicka:2015jba ; multi-singlet SM extensions Goncalves:2025snm ; the Next-to-Minimal Supersymmetric Standard Model Ellwanger:2009dp ; and models featuring scalar or fermionic portals to hidden sectors Arcadi:2024wwg .
+The simplest scalar extension introduces a real, neutral $SU(2)$ singlet with a global $\mathbb{Z}_{2}$ symmetry to ensure stability Burgess et al.
+(2001); Athron and others (2017).
+Alternatively, a minimal fermionic option is a Weyl fermion, commonly referred to as a sterile neutrino Cirelli et al.
+(2024).
+Other proposed models include the Inert Doublet Model, which extends the Higgs sector with a dark doublet Ilnicka et al.
+(2016); multi-singlet SM extensions Gonçalves et al.
+(2025); the Next-to-Minimal Supersymmetric Standard Model Ellwanger et al.
+(2010); and models featuring scalar or fermionic portals to hidden sectors Arcadi et al.
+(2024).
 Most of these models include extra scalar fields, doublet or singlet under the SM gauge group, and often new discrete symmetries in order to stabilize the dark sector.
-In exploring such theories, one must start with the theoretical constraints, such as guaranty of a bounded from below potential, assurance of a global minimum, and perturbative unitarity.
-In any consistent quantum field theory, probability conservation is encoded in the unitarity of the $S$-matrix Schwartz:2014sze ; Logan:2022uus .
-In scalar extensions motivated by dark matter, this requirement becomes particularly relevant: new scalar fields can induce scattering amplitudes that grow with energy, and enforcing unitarity imposes bounds on their masses and couplings.
-For $2\to 2$ scattering processes, unitarity constraints can be applied to the partial-wave decomposition of the amplitudes, with the strongest high-energy limits coming from the zeroth partial-wave amplitude Schwartz:2014sze ; Logan:2022uus .
-In the high-energy limit, the Equivalence Theorem Veltman:1989ud ensures that the scattering of longitudinally polarized gauge bosons can be replaced by the corresponding Goldstone bosons, allowing these scalar amplitudes to capture the dominant contributions to unitarity bounds.
-Historically, Lee, Quigg, and Thacker applied this method to derive an upper bound on the Higgs boson mass, $M_{H}^{2}\leq 8\pi\sqrt{2}/(3G_{F})\approx 1\textrm{TeV}^{2}$ Lee:1977yc ; Lee:1977eg .
-Unitarity has also been applied to the general two-Higgs-doublet model (2HDM) with explicit CP violation Ginzburg:2005dt ; Kanemura:2015ska , to the general $N$-Higgs-doublet model Bento:2017eti , also with fermions Bento:2018fmy , and to constrain large scalar multiplets Hally:2012pu ; Milagre:2024wcg .
+Any consistent exploration of these theories must begin with the enforcement of basic theoretical requirements, such as the boundedness from below of the scalar potential, the existence of a global minimum, and perturbative unitarity, the latter being the main focus of the present work.
+In a consistent quantum field theory, probability conservation is encoded in the unitarity of the $S$-matrix Schwartz (2014); Logan (2022).
+In scalar extensions of the Standard Model, this requirement is particularly restrictive, as the presence of additional scalar degrees of freedom can lead to scattering amplitudes that grow with energy.
+Imposing perturbative unitarity, therefore, yields non-trivial constraints on the masses and couplings of the new scalars, ensuring the validity of perturbation theory.
+For $2\to 2$ scattering processes, perturbative unitarity constraints can be applied to the partial-wave decomposition of the amplitudes, with the strongest high-energy limits coming from the zeroth partial-wave amplitude Schwartz (2014); Logan (2022).
+In the high-energy limit, the equivalence theorem Veltman (1990) ensures that the scattering of longitudinally polarized gauge bosons can be replaced by the corresponding Goldstone bosons, allowing these scalar amplitudes to capture the dominant contributions to perturbative unitarity bounds.
+Historically, Lee, Quigg, and Thacker applied this method to derive an upper bound on the Higgs boson mass, $M_{H}^{2}\leq 8\pi\sqrt{2}/(3G_{F})\approx 1\textrm{TeV}^{2}$ Lee et al.
+(1977a, b).
+Perturbative unitarity has also been applied to the general two-Higgs-doublet model (2HDM) with explicit CP violation Ginzburg and Ivanov (2005); Kanemura and Yagyu (2015), to the general $N$-Higgs-doublet model Bento et al.
+(2017), also with fermions Bento et al.
+(2018), and to constrain large scalar multiplets Hally et al.
+(2012); Milagre and Lavoura (2024).
+Many other examples can be found in the literature, including Llewellyn Smith (1973); Cornwall et al.
+(1974); Weldon (1984); Casalbuoni et al.
+(1986, 1988); Gunion et al.
+(1991); Maalampi et al.
+(1991); Kanemura et al.
+(1993); Akeroyd et al.
+(2000); Horejsi and Kladiva (2006); El Kaffas et al.
+(2007); Grinstein et al.
+(2014); Nagai et al.
+(2015); Englert et al.
+(2016); Cacchio et al.
+(2016); Harris and Logan (2017); Goodsell and Staub (2018, 2019); Bento et al.
+(2023); Urquía-Calderón et al.
+(2025); Blažek et al.
+(2025).
 More recently, perturbative unitarity analyses have been extended beyond the traditional $2\to 2$ limit.
 For instance, Ref.
-Bresciani:2025toe studies unitarity in general $M\to N$ scattering processes with $M,N\geq 2$.
-In particular, in inflationary models featuring a non-minimal coupling with the Ricci scalar, partial-wave unitarity bounds on $2\to N$ processes have improved earlier estimates of the energy scale at which new-physics effective interactions invalidate the perturbative expansion Steingasser:2025txd .
+Bresciani et al.
+(2025) studies unitarity bounds in general $M\to N$ scattering processes with $M,N\geq 2$.
+In particular, in inflationary models featuring a non-minimal coupling with the Ricci scalar, partial-wave unitarity bounds on $2\to N$ processes have improved earlier estimates of the energy scale at which new-physics effective interactions invalidate the perturbative expansion Steingasser et al.
+(2025).
 In this work, we extend the analysis of Ref.
-Bento:2017eti by constructing an $SU(2)\times U(1)$-symmetric model with an arbitrary number of scalar fields (doublets, neutral singlets and charged singlets) and by imposing perturbative unitarity constraints in the high-energy limit.
-Using partial-wave unitarity for tree-level $2\to 2$ scattering, we derive bounds on the quartic couplings.
-In Section II, we present the most general renormalizable $SU(2)\times U(1)$-symmetric model in the high-energy limit, composed of an arbitrary number of $SU(2)$ scalar doublets with hypercharge $\tfrac{1}{2}$ and singlets that are either neutral or carry hypercharge 1.
-We derive constraints on the Lagrangian couplings imposed by hermiticity and identify additional symmetry relations arising from the interchange of field indices.
+Bento et al.
+(2017) by constructing an $SU(2)\times U(1)$-symmetric model with an arbitrary number of scalar fields which may either be singlets or doublets under $SU(2)$.
+The hypercharge assignments are chosen such that the new scalars are either electrically neutral or carry a single unit of electric charge.
+With this framework, we then require all relevant $2\to 2$ scattering processes to satisfy perturbative unitarity at tree-level and in the high-energy limit, thereby deriving upper bounds on the quartic couplings of this class of theories.
+In Section II, we present the most general renormalizable $SU(2)\times U(1)$-symmetric model in the high-energy limit, composed of an arbitrary number of $SU(2)$ scalar doublets with hypercharge $\tfrac{1}{2}$ and singlets that may either be electrically neutral, or carry hypercharge 1.
+We derive relations among the couplings by imposing Hermiticity of the scalar potential and by eliminating redundant gauge-invariant operators.
 In Section III, we construct the scattering matrix from second derivatives of the quartic potential, diagonalize it to obtain eigenamplitudes, and impose perturbative unitarity, $|\Lambda|\leq 8\pi$, to constrain the quartic couplings.
@@
 In Appendix A, we illustrate the benefits of labeling states not only by $Q$ and $Y$, but also by total isospin $T$, using the Standard Model as a case study in Appendix B.
-This classification facilitates the construction of scattering matrices and the identification of independent channels, for which we present a minimal basis (meaning that we discuss the selection of basis states that allows all independent eigenvalues to be determined).
+This classification facilitates the construction of scattering matrices and the identification of independent channels, for which we present a minimal basis (meaning that we discuss the selection of basis states that allow all independent eigenvalues to be determined).
 We provide explicit formulas for the potential and the matrix elements of all scattering processes allowed in our class of models.
-In Section IV, we present the Mathematica notebook BounDS that allows for the automatic calculation of the quartic part of the scalar potential and of all scattering matrices in models of this type; the user must only input the particle content, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries [^1] These can either be discrete or continuous, abelian or non-abelian.
-.
+In Section IV, we present the Mathematica notebook BounDS that allows for the automatic calculation of the quartic part of the scalar potential and of all scattering matrices in models of this type; the user must only input the particle content, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries [^1] These can either be discrete or continuous, Abelian or non-Abelian..
 We make the notebook publicly available at
@@
 In specific models, it is sometimes possible to write the parameters of the potential in terms of scalar masses and/or mixing angles.
-That involves deciding on terms of dimension two and/or three, defining the vacuum of the theory, writing all (neutral and charged) scalar mass matrices, diagonalizing them in order to obtain masses and mixing angles in terms of potential parameters, and then inverting those relations.
-This must be done on a case-by-case basis, and lies beyond the scope of this work.
+That implies defining gauge-invariant terms of dimension two and/or three, defining the vacuum of the theory, writing all (neutral and charged) scalar mass matrices, diagonalizing them to obtain masses and mixing angles in terms of potential parameters, and then inverting those relations.
+This must be done on a case-by-case basis and lies beyond the scope of this work.
 We present our conclusions in Section VI.
@@
 Consider a gauge theory symmetric under the $SU(2)\times U(1)$ group with an arbitrary number of scalar fields (and their charge conjugates) that are either doublets or singlets of $SU(2)$.
-We define the electric charge $Q$ as the sum of the third component of isospin $T_{3}$, and the hypercharge $Y$, and restrict our discussion to models where the electric charge of the scalar particles is either singly charged [^2] In units of the electron charge $e$.
-or neutral.
+We define the electric charge $Q$ as the sum of the third component of isospin $T_{3}$, and the hypercharge $Y$, and restrict our discussion to models where the scalar fields are either electrically neutral or carry a single unit of electric charge.
 As a consequence, this theory may have $n_{D}$ $SU(2)$ scalar doublets with $Y=1/2$ which we denote by
-$\Phi_{i}=\begin{pmatrix}\phi_{i}^{+},\,\phi_{i}^{0}\end{pmatrix}^{T},\quad i=1,\ldots,n_{D},$ (1)
+$\Phi_{i}=\begin{pmatrix}\phi_{i}^{+},\,\phi_{i}^{0}\end{pmatrix}^{T},\quad i=1,\ldots,n_{D};$ (1)
 $n_{c}$ $SU(2)$ complex scalar singlets with $Y=1$ which we denote by
-$\varphi_{i}^{+},\quad i=1,\ldots,n_{c},$ (2)
-and $n_{n}$ $SU(2)$ real scalar singlets with $Y=0$ which we denote by
+$\varphi_{i}^{+},\quad i=1,\ldots,n_{c};$ (2)
+$n_{n}$ $SU(2)$ real scalar singlets with $Y=0$ which we denote by
 $\chi_{i},\quad i=1,\ldots,n_{n}.$ (3)
-Notice that models with $m$ $SU(2)$ complex scalar singlets with $Y=0$ are particular cases of models with $n_{n}=2m$ $SU(2)$ real scalar singlets with $Y=0$; hence, we do not treat them separately.
+Notice that models with $m$ $SU(2)$ complex scalar singlets with $Y=0$ are particular cases of models with $n_{n}=2m$ real scalar singlets with additional $\mathbb{Z}_{2}$ symmetries; hence, we do not treat them separately.
 The oblique radiative corrections for this class of models has been thoroughly studied in Refs.
-Grimus:2007if ; Grimus:2008nb .
+Grimus et al.
+(2008a, b).
 II.2 Scalar potential
-In this work, we are interested in constraining the parameters of such models by imposing partial wave unitarity bounds on $2\to 2$ scattering processes.
-We perform all calculations in the high-energy limit where all contributions from propagators vanish, deeming the quartic interactions in the potential the dominant contributions.
+In this work, we are interested in constraining the parameters of models like the one presented in Section II by imposing partial-wave unitarity bounds on $2\to 2$ scattering processes.
+We perform all calculations in the high-energy regime, where the contributions from propagators in the $s$, $t$, and $u$-channels vanish, deeming the quartic interactions in the potential the dominant contributions.
 The most general renormalizable quartic part of the scalar potential of such a theory can be written as: [^2] For the simplest model including the $\kappa_{ab,cd}$ terms, see Sec.
@@
 Although compact, not all couplings in this notation are independent.
-By swapping fields and requiring the scalar potential to be hermitian, we derive the following relations between quartic couplings:
+By requiring the scalar potential to be Hermitian and rearranging gauge invariant bilinears [^3] By rearranging gauge invariant bilinears, we mean that $\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\left(\Phi_{c}^{\dagger}\Phi_{d}\right)=\left(\Phi_{c}^{\dagger}\Phi_{d}\right)\left(\Phi_{a}^{\dagger}\Phi_{b}\right)\quad\Rightarrow\quad\lambda_{ab,cd}=\lambda_{cd,ab}.$ (5) , we derive the following relations between quartic couplings:
 $\lambda_{ab,cd}$ $=$ $\lambda_{ba,dc}^{*}\ =\ \lambda_{cd,ab},$ (6)
@@
 may have contributions from $s$, $t$, and $u$-channels, and contact interactions.
-However, as we take the limit where the Mandelstam variables $s$, $t$, and $u$ go to infinity, the contributions from $s$, $t$, and $u$-channels vanish, respectively  [^4] For an interesting example see, for example, Appendix A of MiguelMSc ..
-Consequently, in the high-energy limit, only the quartic interactions involving the external scalars contribute to the tree-level amplitude, allowing us to write it as
+However, as we take the limit where the Mandelstam variables $s$, $t$, and $u$ go to infinity, the contributions from $s$, $t$, and $u$-channels vanish, respectively  [^4] For an interesting example, see, for instance, Appendix A of Bento (2017)..
+Consequently, in the high-energy limit, only the quartic interactions involving the external scalars contribute to the tree-level amplitude, allowing us to identify
 $\mathcal{M}\left[A_{a}\,B_{b}\to C_{c}\,D_{d}\right]\ =\ -\frac{\partial^{4}V_{4}}{\partial A_{a}\,\partial B_{b}\,\partial C_{c}^{*}\,\partial D_{d}^{*}}.$ (21)
 Any square-integrable function may be expressed as an expansion in a complete set of basis functions.
-To study perturbative unitarity, it is particularly convenient to choose the basis of Legendre polynomials $P_{J}(\cos\theta)$, where $J$ is the total orbital angular momentum of the final state, and $\theta$ the scattering angle.
-Expanding the amplitude in this basis defines the so-called partial-wave expansion Hally:2012pu ; Milagre:2024wcg :
+To study perturbative unitarity, it is particularly convenient to choose the basis of Legendre polynomials $P_{J}(\cos\theta)$, where $J$ is the total angular momentum of the final state, and $\theta$ the scattering angle.
+Expanding the amplitude in this basis defines the so-called partial-wave expansion Hally et al.
+(2012); Milagre and Lavoura (2024):
 $\mathcal{M}(\cos\theta)=16\pi\sum_{J=0}^{\infty}a_{J}\left(2J+1\right)P_{J}(\cos\theta).$ (22)
@@
 $a_{J}=\frac{2J+1}{32\pi}\int_{-1}^{1}\mathcal{M}(\cos\theta)\,P_{J}(\cos\theta)\ \mathrm{d}\cos\theta,$ (23)
-and are constrained by tree-level partial-wave unitarity as Hally:2012pu ; Milagre:2024wcg
+and are constrained by tree-level partial-wave unitarity through Hally et al. (2012); Milagre and Lavoura (2024)
 […]
-Using the fact that, in the high-energy limit, the amplitude is independent of the scattering angle, and that $P_{0}(\cos\theta)=1$, the most stringent partial-wave unitarity bound arises by applying the second condition of Eq.
-(23) to the zeroth partial wave $a_{0}$.
-Thus, using Eq.
-(20), all quartic couplings must satisfy
+At tree-level, all $a_{J}$ are real, so the three inequalities in Eq.
+(24) reduce to $\left|a_{J}\right|\leq\frac{1}{2}$.
+In the high-energy limit, the scattering amplitudes are independent of the scattering angle, and since $P_{0}(\cos\theta)=1$, the strongest bound arises from the zeroth partial wave, $a_{0}$.
+The perturbative unitarity condition can then be written at the level of the amplitude as
+$\left|a_{0}\right|\ =\ \frac{1}{16\pi}\Big|\mathcal{M}\left[A_{a}\,B_{b}\to C_{c}\,D_{d}\right]\Big|\ \leq\ \frac{1}{2},$ (25)
+or, using Eq. (21), at the level of the quartic couplings as
 $16\pi\left|a_{0}\right|\ =\ \left|N_{ab}N_{cd}\,\frac{\partial^{4}V_{4}}{\partial A_{a}\,\partial B_{b}\,\partial C_{c}^{*}\,\partial D_{d}^{*}}\right|\leq 8\pi,$ (26)
@@
 III.2 Scattering matrices
-One may go a step further by employing the method of coupled-channel analysis Logan:2022uus ; Ginzburg:2005dt ; Kanemura:2015ska .
-This approach takes advantage of the fact that partial wave unitarity bounds may be imposed on any specific process, but also to any superposition of states, provided they have the same quantum numbers.
-By organizing the zeroth partial waves of such scatterings into a coupled-channel matrix, the unitarity condition translates into a bound on its eigenvalues.
-The most stringent constraint is obtained by requiring that the modulus of the largest eigenvalue to remain below $8\pi$.
+One may go a step further by employing the method of coupled-channel analysis Logan (2022); Ginzburg and Ivanov (2005); Kanemura and Yagyu (2015).
+This approach takes advantage of the fact that partial-wave unitarity bounds may be imposed on any specific process, but also to any superposition of states, provided they have the same quantum numbers.
+By organizing the zeroth partial waves of such scatterings into a coupled-channel matrix, the perturbative unitarity condition translates into a bound on its eigenvalues.
+The most stringent constraint is obtained by requiring that the modulus of the largest eigenvalue to remain below $8\pi$, as derived in Eq.
+(26).
 Once again, note that $J=0$ partial waves should be multiplied by a $1/\sqrt{2}$ for every pair of identical particles in the initial or final state.
-We can classify two-particle initial and final states according to their total electric charge $Q$ and hypercharge $Y$, following the approach in Bento:2017eti .
+We can classify two-particle initial and final states according to their total electric charge $Q$ and hypercharge $Y$, following the approach in Bento et al.
+(2017).
 In addition to $Q$ and $Y$, any scattering process involving $SU(2)$ doublets must also conserve total isospin, $T$.
-Following the results of Appendix A, we list in Table 1 the most minimal set of independent two-particle states, labeled by the quantum numbers $|Q,Y,T\rangle$.
+Following the results derived in Appendix A, we list in Table 1 the minimal set of independent two-particle states, labeled by the quantum numbers $|Q,Y,T\rangle$.
 Throughout we use the notation
@@
 $\Phi_{i}\,\Phi_{j}^{*}$ $\equiv$ $\frac{1}{\sqrt{2}}\left(\phi^{+}_{i}\,\phi^{-}_{j}+\phi^{0}_{i}\,\phi^{0*}_{j}\right).$ (29)
+Table 1: Basis of two-particle states labeled by $|Q,Y,T\rangle$. This table had been reduced by including only non-redundant sets of states; see Appendix A for more details.
 $|Q,Y,T\rangle$ State Conditions Dimensionality
@@
 $|0,0,0\rangle$ $\left\{\Phi_{i}\Phi_{j}^{*},\ \varphi_{i}^{+}\varphi_{j}^{-},\ \chi_{i}\chi_{j}\right\}$ $\{\,\textrm{---}\,,\,\textrm{---}\,,\,i\leq j\}$ $n_{n}(n_{n}+1)/2+n_{D}^{2}+n_{c}^{2}$
-Table 1: Basis of two-particle states labeled by $|Q,Y,T\rangle$. This table had been reduced by including only non-redundant sets of states; see Appendix A for more details.
-Looking at Table 1, we notice that the non-redundant sets of states could be classified exclusively in terms of $Y$ and $T$, as is done for the 2HDM in Ginzburg:2005dt .
+Looking at Table 1, we notice that the non-redundant sets of states could be classified exclusively in terms of $Y$ and $T$, as is done for the 2HDM in Ginzburg and Ivanov (2005).
 On the other hand, Ref.
-Bento:2017eti advocated $Q$ and $Y$ when considering all states, since then, as can be seen in Table 9 of Appendix A, there are states with different $Q$ for the same $(Y,T)$.
-We find it best to make the bridge and consider $|Q,Y,T\rangle$.
+Bento et al.
+(2017) advocated for $Q$ and $Y$ when considering all states, since, as can be seen in Table 9 of Appendix A, there are states with different $Q$ for the same $(Y,T)$.
+Nevertheless, when considering scalars in representations of $SU(2)\times U(1)$ other than the ones we use here, the labeling of two-particle states by $Q$, $Y$, and $T$ is the most convenient as it leads to the minimal set of scattering matrices.
+Therefore, we find it best to make the bridge and consider $|Q,Y,T\rangle$.
 Using Eqs. (4)–(19), and (26), the scattering matrix elements for the relevant processes read
@@
 $16\pi\,a_{0}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right$] $=$ $\delta_{ca,db},$ (31)
-$16\pi\,a_{0}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right$] $=$ $N_{ab}N_{cd}\,\left(2\lambda_{ca,db}+2\lambda_{da,cb}\right),$ (30)
-$16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\phi^{+}_{[c}\,\phi^{0}_{d]}\right$] $=$ $\left(2\lambda_{ca,db}-2\lambda_{da,cb}\right),$ (31)
+$16\pi\,a_{0}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right$] $=$ $2N_{ab}N_{cd}\,\left(\lambda_{ca,db}+\lambda_{da,cb}\right),$ (32)
+$16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\phi^{+}_{[c}\,\phi^{0}_{d]}\right$] $=$ $2\left(\lambda_{ca,db}-\lambda_{da,cb}\right),$ (33)
 $16\pi\,a_{0}\left[\phi^{+}_{[a}\,\phi^{0}_{b]}\to\varphi^{+}_{c}\chi_{d}\right]$ $=$ $2\sqrt{2}\,i\,\kappa_{ba,cd},$ (34)
@@
 Note that in Ref.
-Bento:2022vsb a method is proposed based on principal minors, that forgoes diagonalization.
+Bento et al.
+(2022) a method is proposed based on principal minors, that forgoes diagonalization.
 In specific numerical simulations, this is likely to be computationally preferable in all models dealing with large scattering matrices.
@@
 To optimize this process and reduce the burden on the high-energy physics phenomenology community, we have developed the Mathematica tool BounDS that automates the calculation.
-The user simply has to specify the values of $n_{D}$, $n_{c}$, and $n_{n}$, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries.
+The user simply has to specify the values of $n_{D}$, $n_{c}$, and $n_{n}$, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries which may either be discrete or continuous.
 BounDS carries out the necessary steps to:
@@
 The results of the next section are derived from this notebook.
-V Unitarity Bounds for particular cases
-We now proceed to study unitarity bounds in specific scenarios by specifying the values of $n_{D}$, $n_{c}$, and $n_{n}$, as well as any additional symmetries.
+V Perturbative Unitarity Bounds for particular cases
+We now proceed to study perturbative unitarity bounds in specific scenarios by specifying the values of $n_{D}$, $n_{c}$, and $n_{n}$, as well as any additional symmetries.
 The cases studied explicitly in this work are summarized in Table 2.
+Table 2: List of models presented as examples in this article.
 $_{D} nn$$_{n}$ $n_{c}$ Symmetries Ref.
-1 0 0 Lee:1977eg
+1 0 0 Lee et al. (1977b)
 1 1 0
-2 0 0 $\mathbb{Z}_{2}$ Ginzburg:2005dt
-2 0 0 Ginzburg:2005dt
-1 2 0 Muhlleitner:2020wwk generalized
+2 0 0 $\mathbb{Z}_{2}$ Ginzburg and Ivanov (2005)
+2 0 0 Ginzburg and Ivanov (2005)
+1 2 0 Mühlleitner et al. (2022) generalized
 2 1 0 $\mathbb{Z}_{2}$
-2 2 0 $\mathbb{Z}_{2}\otimes\mathbb{Z}^{\prime}_{2}$ Boto:2025mmn
+2 2 0 $\mathbb{Z}_{2}\otimes\mathbb{Z}^{\prime}_{2}$ Boto et al. (2025)
 2 1 1
-3 0 0 $\mathbb{Z}_{3}$ Bento:2017eti
-Table 2: Models presented as examples in this article.
+3 0 0 $\mathbb{Z}_{3}$ Bento et al. (2017)
 V.1 The Standard Model
@@
 $M_{|0,0,0\rangle}=6\lambda_{11,11}.$ (48)
-V.1.2 Unitarity Bounds
-Applying the partial wave unitarity condition of Eq. (24) to the eigenvalues of these scattering matrices, we find
+V.1.2 Perturbative Unitarity Bounds
+Applying the partial-wave unitarity condition of Eq. (26) to the eigenvalues of these scattering matrices, immediately leads to
 $|2\lambda_{11,11}|\leq 8\pi,\quad|6\lambda_{11,11}|\leq 8\pi\quad\Longrightarrow\quad\lambda_{11,11}\leq\frac{4\pi}{3}.$ (49)
 In order to facilitate the comparison of our results with those of Ref.
-Bento:2017eti , it is useful to provide a short dictionary of notations.
+Bento et al.
+(2017), it is useful to provide a short dictionary of notations.
 Table 3 lists the correspondence between the coupling employed in this work and the one used in Ref.
-Bento:2017eti .
-Term Our Notation Notation in Bento:2017eti
+Bento et al.
+(2017).
+Table 3: Comparison of coupling notation.
+Term Our Notation Notation in Bento et al. (2017)
 $(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\lambda$
-Table 3: Comparison of coupling notation.
-This recovers the classic results of Lee:1977eg .
+This recovers the classic results of Lee et al. (1977b).
 V.2 1 Scalar Doublet and 1 Neutral Scalar Singlet
@@
 $M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=2\gamma_{11,11},$ (52)
-$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&2\sqrt{2}\gamma_{11,11}\\ 2\gamma_{11,11}&12\beta_{11,11}\\ \end{bmatrix}$. (51)
-V.2.2 Unitarity Bounds
-We apply the partial wave unitarity condition of Eq. (24) to each of the eigenvalues of the zero partial wave amplitude matrix and find
+$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&2\gamma_{11,11}\\ 2\gamma_{11,11}&12\beta_{11,11}\\ \end{bmatrix}$. (53)
+V.2.2 Perturbative Unitarity Bounds
+We apply the partial-wave unitarity condition of Eq. (26) to each of the eigenvalues of the zero partial-wave amplitude matrix and find
 $\left|6\beta_{11,11}+3\lambda_{11,11}\pm\sqrt{9\left(\lambda_{11,11}-2\beta_{11,11}\right){}^{2}+4\gamma_{11,11}^{2}}\right|$ $\leq$ $8\pi,$ (54)
@@
 (60) we have employed the standard notation of Ref.
-Ginzburg:2005dt and, without loss of generality, choose $\lambda_{5}$ to be real.
+Ginzburg and Ivanov (2005) and, without loss of generality, choose $\lambda_{5}$ to be real.
 The mapping between the two notations is given in Table 4.
-Term Our Notation Notation in Ginzburg:2005dt
+Table 4: Comparison of coupling notation.
+Term Our Notation Notation in Ginzburg and Ivanov (2005)
 $(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\frac{\lambda_{1}}{2}$
@@
 $(\Phi_{2}^{\dagger}\Phi_{1})^{2}$ $\lambda_{12,12}^{*}$ $\frac{\lambda_{5}}{2}$
-Table 4: Comparison of coupling notation.
 V.3.1 Scattering Matrices
@@
 V.3.2 Eigenvalues
-By imposing the unitarity condition of Eq. (24) to each of the eigenvalues of the zero partial wave amplitude matrix we derive
+By imposing the unitarity condition of Eq. (26) to each of the eigenvalues of the zero partial-wave amplitude matrix we derive
 $\left|\lambda_{3}\pm\lambda_{4}\right|$ $\leq$ $8\pi,$ (65)
@@
 $\frac{1}{2}\left|3\lambda_{1}+3\lambda_{2}\pm\sqrt{9\left(\lambda_{1}-\lambda_{2}\right)^{2}+4\left(2\lambda_{3}+\lambda_{4}\right)^{2}}\right|$ $\leq$ $8\pi.$ (70)
-We find that our results are in agreement with those derived in Ginzburg:2005dt .
+We find that our results are in agreement with those derived in Ginzburg and Ivanov (2005).
 V.4 The general 2HDM
@@
 $\left[\frac{\lambda_{5}}{2}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{6}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{1}^{\dagger}\Phi_{2}\right)+\lambda_{7}\left(\Phi_{1}^{\dagger}\Phi_{2}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)+\text{h.c.}\right],$
-where, in Eq. (70) we have employed, once again, the standard notation of Ref. Ginzburg:2005dt . The mapping between the two notations is given in Table 5.
-Term Our Notation Notation in Ginzburg:2005dt
+where, in Eq. (72) we have employed, once again, the standard notation of Ref. Ginzburg and Ivanov (2005). The mapping between the two notations is given in Table 5.
+Table 5: Comparison of coupling notation.
+Term Our Notation Notation in Ginzburg and Ivanov (2005)
 $(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\frac{\lambda_{1}}{2}$
@@
 $(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{2})$ $2\lambda_{12,22}$ $\lambda_{7}$
-Table 5: Comparison of coupling notation.
 V.4.1 Scattering Matrices
 The set of non-zero scattering matrices for this model reads
-$M_{|2,1,1\rangle}=\begin{bmatrix}\lambda_{1}&\sqrt{2}\,\lambda_{6}^{*}&\lambda_{5}^{*}\\ \sqrt{2}\,\lambda_{6}&\lambda_{3}+\lambda_{4}&\sqrt{2}\,\lambda_{7}^{*}\\ \lambda_{5}&\sqrt{2}\,\lambda_{7}&\lambda_{2}\end{bmatrix}$, (71)
-$M_{|1,0,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{6}&\lambda_{6}^{*}&\lambda_{4}\\ \lambda_{6}^{*}&\lambda_{3}&\lambda_{5}^{*}&\lambda_{7}^{*}\\ \lambda_{6}&\lambda_{5}&\lambda_{3}&\lambda_{7}\\ \lambda_{4}&\lambda_{7}&\lambda_{7}^{*}&\lambda_{2}\end{bmatrix}$, (72)
+$M_{|2,1,1\rangle}=\begin{bmatrix}\lambda_{1}&\sqrt{2}\,\lambda_{6}&\lambda_{5}\\ \sqrt{2}\,\lambda_{6}^{*}&\lambda_{3}+\lambda_{4}&\sqrt{2}\,\lambda_{7}\\ \lambda_{5}^{*}&\sqrt{2}\,\lambda_{7}^{*}&\lambda_{2}\end{bmatrix}$, (73)
+$M_{|1,0,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{6}^{*}&\lambda_{6}&\lambda_{4}\\ \lambda_{6}&\lambda_{3}&\lambda_{5}&\lambda_{7}\\ \lambda_{6}^{*}&\lambda_{5}^{*}&\lambda_{3}&\lambda_{7}^{*}\\ \lambda_{4}&\lambda_{7}^{*}&\lambda_{7}&\lambda_{2}\end{bmatrix}$, (74)
 $M_{|0,0,0\rangle}=\begin{bmatrix}3\lambda_{1}&3\lambda_{6}^{*}&3\lambda_{6}&2\lambda_{3}+\lambda_{4}\\ 3\lambda_{6}&\lambda_{3}+2\lambda_{4}&3\lambda_{5}&3\lambda_{7}\\ 3\lambda_{6}^{*}&3\lambda_{5}^{*}&\lambda_{3}+2\lambda_{4}&3\lambda_{7}^{*}\\ 2\lambda_{3}+\lambda_{4}&3\lambda_{7}^{*}&3\lambda_{7}&3\lambda_{2}\end{bmatrix}$, (75)
 $M_{|1,1,0\rangle}=\lambda_{3}-\lambda_{4}.$ (76)
-Our scattering matrices agree with those in Ginzburg:2005dt .
+Our scattering matrices agree with those in Ginzburg and Ivanov (2005).
 If we set $\lambda_{6}=\lambda_{7}=0$ (along with their complex conjugates), we re-obtain, as expected, the results presented in Section V.3.1.
-V.4.2 Unitarity Bounds
+V.4.2 Perturbative Unitarity Bounds
 The eigenvalues of the scattering matrices in this case are, in general, too complex to write down in closed form.
-However, some can be computed analytically, and unitarity bounds are imposed accordingly,
+However, some can be computed analytically, and perturbative unitarity bounds are imposed accordingly,
 $|\lambda_{3}-\lambda_{4}|\leq 8\pi.$ (77)
@@
 $M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&2\gamma_{11,12}\\ 2\gamma_{11,12}&2\gamma_{11,22}\end{bmatrix}$, (80)
-$M_{|0,0,0\rangle}=\begin{bmatrix}6\,\lambda_{11,11}&2\sqrt{2}\,\gamma_{11,11}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{11,22}\\ 2\,\gamma_{11,11}&12\,\beta_{11,11}&12\sqrt{2}\,\beta_{11,12}&12\,\beta_{11,22}\\ 2\sqrt{2}\,\gamma_{11,12}&12\sqrt{2}\,\beta_{11,12}&24\,\beta_{11,22}&12\sqrt{2}\,\beta_{12,22}\\ 2\,\gamma_{11,22}&12\,\beta_{11,22}&12\sqrt{2}\,\beta_{12,22}&12\,\beta_{22,22}\end{bmatrix}$.
-(79)
-V.5.2 Unitarity Bounds
+$M_{|0,0,0\rangle}=\left[\begin{array}[]{cccc}6\lambda_{11,11}&2\gamma_{11,11}&2\sqrt{2}\gamma_{11,12}&2\gamma_{11,22}\\ 2\gamma_{11,11}&12\beta_{11,11}&12\sqrt{2}\beta_{11,12}&12\beta_{11,22}\\ 2\sqrt{2}\gamma_{11,12}&12\sqrt{2}\beta_{11,12}&24\beta_{11,22}&12\sqrt{2}\beta_{12,22}\\ 2\gamma_{11,22}&12\beta_{11,22}&12\sqrt{2}\beta_{12,22}&12\beta_{22,22}\\ \end{array}\right].$ (81)
+V.5.2 Perturbative Unitarity Bounds
 Again, certain eigenvalues of the scattering matrices are too complicated to be expressed in closed form.
-However, others can be written analytically, and, for those, the corresponding unitarity bounds are then applied, as:
+However, others can be written analytically, and, for those, the corresponding perturbative unitarity bounds are then applied, as:
 $\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{4\gamma_{11,12}^{2}+\left(\gamma_{11,11}-\gamma_{11,22}\right){}^{2}}\right|$ $\leq 8\pi,$ (82)
@@
 In order to compare our results with previous work, we consider the model studied in Ref.
-Muhlleitner:2020wwk , where the SM is extended by a neutral complex scalar singlet.
+Mühlleitner et al.
+(2022), where the SM is extended by a neutral complex scalar singlet.
 In their notation, the quartic part of the scalar potential reads
 $V_{4}=\frac{\lambda}{4}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\frac{\delta_{2}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)|S|^{2}+\frac{d_{2}}{2}|S|^{4},$ (85)
-where  [^5] Equivalently, one could choose $S=\frac{\chi_{2}+i\chi_{1}}{\sqrt{2}}$.
-$S=\frac{\chi_{1}+i\chi_{2}}{\sqrt{2}}$.
+where $S=\frac{\chi_{1}+i\chi_{2}}{\sqrt{2}}$.
 Notice that the case with two real singlets in Eq.
@@
 (85).
-But, since the complex scalar singlet can be decomposed into two real scalar singlet fields, $\chi_{1}$ and $\chi_{2}$, we find that our results reproduce those of Muhlleitner:2020wwk , when we set $\gamma_{11,12}=\beta_{11,12}=\beta_{12,22}=0$.
+But, since the complex scalar singlet can be decomposed into two real scalar singlet fields, $\chi_{1}$ and $\chi_{2}$, we find that our results reproduce those of Mühlleitner et al.
+(2022), when we set $\gamma_{11,12}=\beta_{11,12}=\beta_{12,22}=0$, i.e.
+when we impose an additional $\mathbb{Z}_{2}$ that acts as $\chi_{2}\to-\chi_{2}$.
+[^5] Equivalently, one could choose $S=\frac{\chi_{2}+i\chi_{1}}{\sqrt{2}}$ and impose the $\mathbb{Z}_{2}$ symmetry on $\chi_{1}$.
+This procedure is generalizable to models with any number of complex neutral singlet scalars.
 By matching the terms in Eq.
@@
 (78), we derive the mapping between the two notations that is provided in Table 6.
-Term Our Notation Notation in Muhlleitner:2020wwk
+Table 6: Comparison of coupling notation.
+Term Our Notation Notation in Mühlleitner et al. (2022)
 $(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\frac{\lambda}{4}$
@@
 $(\Phi_{1}^{\dagger}\Phi_{1})\chi_{1}\chi_{2}$ $2\gamma_{11,12}$ 0
-Table 6: Comparison of coupling notation.
 V.6 2 Scalar Doublets and 1 Neutral Scalar Singlet with a $\mathbb{Z}_{2}$ symmetry
@@
 $M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&0\\ 0&2\gamma_{22,11}\end{bmatrix}$, (88)
-$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda^{*}_{12,12}&0\\ 2\lambda_{12,12}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (87)
-$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda^{*}_{12,12}\\ 0&0&2\lambda_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (88)
-$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&4\lambda_{11,22}+2\lambda_{12,21}&2\sqrt{2}\,\gamma_{11,11}&0&0\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda_{22,22}&2\sqrt{2}\,\gamma_{22,11}&0&0\\ 2\gamma_{11,11}&2\gamma_{22,11}&12\beta_{11,11}&0&0\\ 0&0&0&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,12}\\ 0&0&0&6\lambda^{*}_{12,12}&2\lambda_{11,22}+4\lambda_{12,21}\end{bmatrix}$, (89)
+$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,12}&0\\ 2\lambda^{*}_{12,12}&2\lambda_{22,22}&0\\ 0&0&2(\lambda_{11,22}+\lambda_{12,21})\end{bmatrix}$, (89)
+$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda_{12,12}\\ 0&0&2\lambda^{*}_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (90)
+$M_{|0,0,0\rangle}=\left[\begin{array}[]{ccccc}6\lambda_{11,11}&2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&2\gamma_{11,11}&0&0\\ 2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&6\lambda_{22,22}&2\gamma_{22,11}&0&0\\ 2\gamma_{11,11}&2\gamma_{22,11}&12\beta_{11,11}&0&0\\ 0&0&0&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,12}\\ 0&0&0&6\lambda_{12,12}^{*}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)\\ \end{array}\right],$ (91)
 $M_{|1,1,0\rangle}=2\left(\lambda_{11,22}-\lambda_{12,21}\right).$ (92)
-V.6.2 Unitarity Bounds
+V.6.2 Perturbative Unitarity Bounds
 Some eigenvalues are too complicated to be written in closed form.
@@
 $\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4\lambda_{12,21}^{2}}\right|$ $\leq 8\pi,$ (97)
-$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq 8\pi.$ (96)
+$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq 8\pi,$ (98)
+$2\left|\lambda_{11,22}+2\lambda_{12,21}\pm 3|\lambda_{12,12}|^{2}\right|$ $\leq 8\pi.$ (99)
 V.7 2 Scalar Doublets with a $\mathbb{Z}_{2}$ symmetry + 2 Neutral Scalar Singlets with a $\mathbb{Z^{\prime}}_{2}$ symmetry
@@
 The set of non-zero scattering matrices for this model reads
-$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&2\gamma_{11,12}&0&0\\ 2\gamma_{11,12}&2\gamma_{11,22}&0&0\\ 0&0&2\gamma_{22,11}&2\gamma_{22,12}\\ 0&0&2\gamma_{22,12}&2\gamma_{22,22}\end{bmatrix}$, (100)
-$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda^{*}_{12,12}&0\\ 2\lambda_{12,12}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (101)
-$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda^{*}_{12,12}\\ 0&0&2\lambda_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (102)
+$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\left[\begin{array}[]{cccc}2\gamma_{22,11}&2\gamma_{22,12}&0&0\\ 2\gamma_{22,12}&2\gamma_{22,22}&0&0\\ 0&0&2\gamma_{11,22}&2\gamma_{11,12}\\ 0&0&2\gamma_{11,12}&2\gamma_{11,11}\\ \end{array}\right],$ (103)
+$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,12}&0\\ 2\lambda_{12,12}^{*}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (104)
+$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda_{12,12}\\ 0&0&2\lambda^{*}_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (105)
 $M_{|1,1,0\rangle}=2\left(\lambda_{11,22}-\lambda_{12,21}\right),$ (106)
@@
 (107) are given, respectively, by
-$A=\begin{bmatrix}24\,\beta_{11,22}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{22,12}&12\sqrt{2}\,\beta_{11,12}&12\sqrt{2}\,\beta_{12,22}\\ 2\sqrt{2}\,\gamma_{11,12}&6\,\lambda_{11,11}&4\,\lambda_{11,22}+2\,\lambda_{12,21}&2\sqrt{2}\,\gamma_{11,11}&2\sqrt{2}\,\gamma_{11,22}\\ 2\sqrt{2}\,\gamma_{22,12}&4\,\lambda_{11,22}+2\,\lambda_{12,21}&6\,\lambda_{22,22}&2\sqrt{2}\,\gamma_{22,11}&2\sqrt{2}\,\gamma_{22,22}\\ 12\sqrt{2}\,\beta_{11,12}&2\,\gamma_{11,11}&2\,\gamma_{22,11}&12\,\beta_{11,11}&12\,\beta_{11,22}\\ 12\sqrt{2}\,\beta_{12,22}&2\,\gamma_{11,22}&2\,\gamma_{22,22}&12\,\beta_{11,22}&12\,\beta_{22,22}\end{bmatrix}$, (105)
-$B=\begin{bmatrix}2\,\lambda_{11,22}+4\,\lambda_{12,21}&6\,\lambda_{12,12}\\ 6\,\lambda_{12,12}^{*}&2\,\lambda_{11,22}+4\,\lambda_{12,21}\end{bmatrix}$. (106)
-V.7.2 Unitarity Bounds
+$A=\begin{bmatrix}24\,\beta_{11,22}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{22,12}&12\sqrt{2}\,\beta_{11,12}&12\sqrt{2}\,\beta_{12,22}\\ 2\sqrt{2}\,\gamma_{11,12}&6\,\lambda_{11,11}&2(2\,\lambda_{11,22}+\,\lambda_{12,21})&2\,\gamma_{11,11}&2\,\gamma_{11,22}\\ 2\sqrt{2}\,\gamma_{22,12}&2(2\,\lambda_{11,22}+\,\lambda_{12,21})&6\,\lambda_{22,22}&2\,\gamma_{22,11}&2\,\gamma_{22,22}\\ 12\sqrt{2}\,\beta_{11,12}&2\,\gamma_{11,11}&2\,\gamma_{22,11}&12\,\beta_{11,11}&12\,\beta_{11,22}\\ 12\sqrt{2}\,\beta_{12,22}&2\,\gamma_{11,22}&2\,\gamma_{22,22}&12\,\beta_{11,22}&12\,\beta_{22,22}\end{bmatrix}$, (108)
+$B=\begin{bmatrix}2(\,\lambda_{11,22}+2\,\lambda_{12,21})&6\,\lambda_{12,12}\\ 6\,\lambda_{12,12}^{*}&2(\,\lambda_{11,22}+2\,\lambda_{12,21})\end{bmatrix}$. (109)
+V.7.2 Perturbative Unitarity Bounds
 As several eigenvalues are too involved to express analytically, we show only the results of imposing partial-wave unitarity on the remaining eigenvalues of the $s$-wave amplitude matrix, that can be computed explicitly,
@@
 $\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq$ $8\pi,$ (114)
-$\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{\gamma_{11,11}^{2}+4\gamma_{11,12}^{2}-2\gamma_{11,11}\gamma_{11,22}+\gamma_{11,22}^{2}}\right|$ $\leq$ $8\pi,$ (112)
-$\left|\gamma_{22,11}+\gamma_{22,22}\pm\sqrt{\gamma_{22,11}^{2}+4\gamma_{22,12}^{2}-2\gamma_{22,11}\gamma_{22,22}+\gamma_{22,22}^{2}}\right|$ $\leq$ $8\pi.$ (113)
-Term Our Notation Notation of Boto:2025mmn
+$\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{4\gamma_{11,12}^{2}+\left(\gamma_{11,11}-\gamma_{11,22}\right){}^{2}}\right|$ $\leq$ $8\pi,$ (115)
+$\left|\gamma_{22,11}+\gamma_{22,22}\pm\sqrt{4\gamma_{22,12}^{2}+\left(\gamma_{22,11}-\gamma_{22,22}\right){}^{2}}\right|$ $\leq$ $8\pi.$ (116)
+Table 7: Comparison of coupling notation.
+Term Our Notation Notation of Boto et al. (2025)
 $(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $\tfrac{\lambda_{1}}{2}$
@@
 $(\Phi_{2}^{\dagger}\Phi_{2})\chi_{1}\chi_{2}$ $2\gamma_{22,12}$ $\tfrac{\lambda_{16}}{2}$
-Table 7: Comparison of coupling notation.
-We verified our results against those reported in Boto:2025mmn and found them to be in agreement.
-Table 7 summarizes the correspondence between the couplings used in our work and those of Boto:2025mmn .
+We verified our results against those reported in Boto et al.
+(2025) and found them to be in agreement.
+Table 7 summarizes the correspondence between the couplings used in our work and those of Boto et al.
+(2025).
 V.8 2 Scalar Doublets + 1 Neutral Scalar Singlet + 1 Charged Scalar Singlet
@@
 $M_{|2,2,0\rangle}=2\alpha_{11,11},$ (118)
-$M_{|2,\frac{3}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}^{*}\\ \delta_{12,11}&\delta_{22,11}\end{bmatrix}$, (116)
-$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}&0&-2i\kappa^{*}_{12,11}\\ \delta^{*}_{12,11}&\delta_{22,11}&2i\kappa^{*}_{12,11}&0\\ 0&-2i\kappa_{12,11}&2\gamma_{11,11}&2\gamma^{*}_{12,11}\\ 2i\kappa_{12,11}&0&2\gamma_{12,11}&2\gamma_{22,11}\end{bmatrix}$, (117)
-$M_{|1,1,0\rangle}=\begin{bmatrix}2\lambda_{11,22}-2\lambda_{12,21}&2i\sqrt{2}\kappa^{*}_{12,11}\\ -2i\sqrt{2}\kappa^{*}_{12,11}&2\zeta_{11,11}\end{bmatrix}$, (118)
-$M_{|1,1,2\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{11,12}^{*}&2\lambda_{12,12}^{*}\\ 2\sqrt{2}\lambda_{11,12}&2\lambda_{11,22}+2\lambda_{12,21}&2\sqrt{2}\lambda_{12,22}^{*}\\ 2\lambda_{12,12}&2\sqrt{2}\lambda_{12,22}&2\lambda_{22,22}\end{bmatrix}$, (119)
-$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{11,12}&2\lambda_{11,12}^{*}&2\lambda_{12,21}\\ 2\lambda_{11,12}^{*}&2\lambda_{11,22}&2\lambda_{12,12}^{*}&2\lambda_{12,22}^{*}\\ 2\lambda_{11,12}&2\lambda_{12,12}&2\lambda_{11,22}&2\lambda_{12,22}\\ 2\lambda_{12,21}&2\lambda_{12,22}&2\lambda_{12,22}^{*}&2\lambda_{22,22}\end{bmatrix}$, (120)
-$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&6\lambda^{*}_{11,12}&6\lambda_{11,12}&4\lambda_{11,22}+2\lambda_{12,21}&\sqrt{2}\delta_{11,11}&2\sqrt{2}\gamma_{11,11}\\ 6\lambda_{11,12}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,12}&6\lambda_{12,22}&\sqrt{2}\delta_{12,11}&2\sqrt{2}\gamma_{12,11}\\ 6\lambda^{*}_{11,12}&6\lambda^{*}_{12,12}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda^{*}_{12,22}&\sqrt{2}\delta^{*}_{12,11}&2\sqrt{2}\gamma^{*}_{12,11}\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda^{*}_{12,22}&6\lambda_{12,22}&6\lambda_{22,22}&\sqrt{2}\delta_{22,11}&2\sqrt{2}\gamma_{22,11}\\ \sqrt{2}\delta_{11,11}&\sqrt{2}\delta^{*}_{12,11}&\sqrt{2}\delta_{12,11}&\sqrt{2}\delta_{22,11}&4\alpha_{11,11}&\sqrt{2}\zeta_{11,11}\\ 2\gamma_{11,11}&2\gamma^{*}_{12,11}&2\gamma_{12,11}&2\gamma_{22,11}&\sqrt{2}\zeta_{11,11}&12\beta_{11,11}\end{bmatrix}$.
-(121)
-V.8.2 Unitarity Bounds
-Since certain eigenvalues are too complex to evaluate analytically, unitarity constraints are shown only for the remaining eigenvalues of the zero partial-wave amplitude matrix. These read
-$\left|\zeta_{11,11}+\lambda_{11,22}-\lambda_{12,21}\pm\sqrt{\left(-\zeta_{11,11}-\lambda_{11,22}+\lambda_{12,21}\right)^{2}-4\left(\zeta_{11,11}\lambda_{11,22}-\zeta_{11,11}\lambda_{12,21}-2|\kappa_{12,11}|^{2}\right)}\right|$ $\leq 8\pi,$ (122)
+$M_{|2,\frac{3}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}\\ \delta_{12,11}^{*}&\delta_{22,11}\end{bmatrix}$, (119)
+$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\left[\begin{array}[]{cccc}2\gamma_{11,11}&2\gamma_{12,11}&0&2i\kappa_{12,11}^{*}\\ 2\gamma_{12,11}^{*}&2\gamma_{22,11}&-2i\kappa_{12,11}^{*}&0\\ 0&2i\kappa_{12,11}&\delta_{11,11}&\delta_{12,11}^{*}\\ -2i\kappa_{12,11}&0&\delta_{12,11}&\delta_{22,11}\\ \end{array}\right],$ (120)
+$M_{|1,1,0\rangle}=\begin{bmatrix}2(\lambda_{11,22}-\lambda_{12,21})&2i\sqrt{2}\kappa^{*}_{12,11}\\ -2i\sqrt{2}\kappa_{12,11}&2\zeta_{11,11}\end{bmatrix}$, (121)
+$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{11,12}&2\lambda_{12,12}\\ 2\sqrt{2}\lambda_{11,12}^{*}&2(\lambda_{11,22}+\lambda_{12,21})&2\sqrt{2}\lambda_{12,22}\\ 2\lambda_{12,12}^{*}&2\sqrt{2}\lambda_{12,22}^{*}&2\lambda_{22,22}\end{bmatrix}$, (122)
+$M_{|1,0,1\rangle}=\left[\begin{array}[]{cccc}2\lambda_{11,11}&2\lambda_{11,12}^{*}&2\lambda_{11,12}&2\lambda_{12,21}\\ 2\lambda_{11,12}&2\lambda_{11,22}&2\lambda_{12,12}&2\lambda_{12,22}\\ 2\lambda_{11,12}^{*}&2\lambda_{12,12}^{*}&2\lambda_{11,22}&2\lambda_{12,22}^{*}\\ 2\lambda_{12,21}&2\lambda_{12,22}^{*}&2\lambda_{12,22}&2\lambda_{22,22}\\ \end{array}\right],$ (123)
+$M_{|0,0,0\rangle}=\left[\begin{array}[]{cccccc}6\lambda_{11,11}&6\lambda_{11,12}^{*}&6\lambda_{11,12}&2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&\sqrt{2}\delta_{11,11}&2\gamma_{11,11}\\ 6\lambda_{11,12}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,12}&6\lambda_{12,22}&\sqrt{2}\delta_{12,11}&2\gamma_{12,11}\\ 6\lambda_{11,12}^{*}&6\lambda_{12,12}^{*}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,22}^{*}&\sqrt{2}\delta_{12,11}^{*}&2\gamma_{12,11}^{*}\\ 2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&6\lambda_{12,22}^{*}&6\lambda_{12,22}&6\lambda_{22,22}&\sqrt{2}\delta_{22,11}&2\gamma_{22,11}\\ \sqrt{2}\delta_{11,11}&\sqrt{2}\delta_{12,11}^{*}&\sqrt{2}\delta_{12,11}&\sqrt{2}\delta_{22,11}&4\alpha_{11,11}&\sqrt{2}\zeta_{11,11}\\ 2\gamma_{11,11}&2\gamma_{12,11}^{*}&2\gamma_{12,11}&2\gamma_{22,11}&\sqrt{2}\zeta_{11,11}&12\beta_{11,11}\\ \end{array}\right].$ (124)
+V.8.2 Perturbative Unitarity Bounds
+Since certain eigenvalues are too complex to evaluate analytically, the perturbative unitarity constraints are shown only for the remaining eigenvalues of the zero partial-wave amplitude matrix.
+These read
+$\left|\zeta_{11,11}+\lambda_{11,22}-\lambda_{12,21}\pm\sqrt{\left(-\zeta_{11,11}-\lambda_{11,22}+\lambda_{12,21}\right)^{2}+8|\kappa_{12,11}|^{2}}\right|$ $\leq 8\pi,$ (125)
 $\frac{1}{2}\left|\delta_{11,11}+\delta_{22,11}\pm\sqrt{\left(\delta_{11,11}-\delta_{22,11}\right)^{2}+4|\delta_{12,11}|^{2}}\right|$ $\leq 8\pi,$ (126)
@@
 where
-$A$ $=$ $\begin{bmatrix}2\lambda_{11,33}+2\lambda_{13,31}&2\sqrt{2}\lambda^{*}_{12,32}\\[2.84526pt] 2\sqrt{2}\lambda_{12,32}&2\lambda_{22,22}\end{bmatrix}$, (128)
-$B$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda^{*}_{12,13}\\[2.84526pt] 2\sqrt{2}\lambda_{12,13}&2\lambda_{22,33}+2\lambda_{23,32}\end{bmatrix}$, (129)
-$C$ $=$ $\begin{bmatrix}2\lambda_{33,33}&2\sqrt{2}\lambda_{13,23}\\[2.84526pt] 2\sqrt{2}\lambda^{*}_{13,23}&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (130)
+$A$ $=$ $\begin{bmatrix}2(\lambda_{11,33}+\lambda_{13,31})&2\sqrt{2}\lambda_{12,32}\\[2.84526pt] 2\sqrt{2}\lambda_{12,32}^{*}&2\lambda_{22,22}\end{bmatrix}$, (131)
+$B$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{12,13}\\[2.84526pt] 2\sqrt{2}\lambda_{12,13}^{*}&2(\lambda_{22,33}+\lambda_{23,32})\end{bmatrix}$, (132)
+$C$ $=$ $\begin{bmatrix}2\lambda_{33,33}&2\sqrt{2}\lambda_{13,23}^{*}\\[2.84526pt] 2\sqrt{2}\lambda_{13,23}&2(\lambda_{11,22}+\lambda_{12,21})\end{bmatrix}$, (133)
 $M_{|1,0,1\rangle}=\textrm{blkdiag}(D,E,F),$ (134)
 where
-$D$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{13,23}&2\lambda^{*}_{12,32}\\ 2\lambda^{*}_{13,23}&2\lambda_{11,33}&2\lambda^{*}_{12,13}\\ 2\lambda_{12,32}&2\lambda_{12,13}&2\lambda_{11,22}\end{bmatrix}$, (132)
+$D$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{13,23}^{*}&2\lambda_{12,32}\\ 2\lambda_{13,23}&2\lambda_{11,33}&2\lambda_{12,13}\\ 2\lambda_{12,32}^{*}&2\lambda_{12,13}^{*}&2\lambda_{11,22}\end{bmatrix}$, (135)
 $E$ $=$ $\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&2\lambda_{13,31}\\ 2\lambda_{12,21}&2\lambda_{22,22}&2\lambda_{23,32}\\ 2\lambda_{13,31}&2\lambda_{23,32}&2\lambda_{33,33}\end{bmatrix}$, (136)
-$F$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{12,32}&2\lambda^{*}_{13,23}\\ 2\lambda^{*}_{12,32}&2\lambda_{11,22}&2\lambda^{*}_{12,13}\\ 2\lambda_{13,23}&2\lambda_{12,13}&2\lambda_{11,33}\end{bmatrix}$, (134)
+$F$ $=$ $\begin{bmatrix}2\lambda_{22,33}&2\lambda_{12,32}^{*}&2\lambda_{13,23}\\ 2\lambda_{12,32}&2\lambda_{11,22}&2\lambda_{12,13}\\ 2\lambda_{13,23}^{*}&2\lambda_{12,13}^{*}&2\lambda_{11,33}\end{bmatrix}$, (137)
 $M_{|0,0,0\rangle}=\textrm{blkdiag}(G,H,I),$ (138)
 where
-$G$ $=$ $\begin{bmatrix}2\lambda_{22,33}+4\lambda_{23,32}&6\lambda^{*}_{13,23}&6\lambda_{12,32}\\ 6\lambda_{13,23}&2\lambda_{11,33}+4\lambda_{13,31}&6\lambda_{12,13}\\ 6\lambda^{*}_{12,32}&6\lambda^{*}_{12,13}&2\lambda_{11,22}+4\lambda_{12,21}\end{bmatrix}$, (136)
-$H$ $=$ $\begin{bmatrix}6\lambda_{11,11}&4\lambda_{11,22}+2\lambda_{12,21}&4\lambda_{11,33}+2\lambda_{13,31}\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda_{22,22}&4\lambda_{22,33}+2\lambda_{23,32}\\ 4\lambda_{11,33}+2\lambda_{13,31}&4\lambda_{22,33}+2\lambda_{23,32}&6\lambda_{33,33}\end{bmatrix}$, (137)
-$I$ $=$ $\begin{bmatrix}2\lambda_{22,33}+4\lambda_{23,32}&6\lambda^{*}_{12,32}&6\lambda_{13,23}\\ 6\lambda_{12,32}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,13}\\ 6\lambda^{*}_{13,23}&6\lambda^{*}_{12,13}&2\lambda_{11,33}+4\lambda_{13,31}\end{bmatrix}$, (138)
+$G$ $=$ $\begin{bmatrix}2(\lambda_{22,33}+2\lambda_{23,32})&6\lambda^{*}_{13,23}&6\lambda_{12,32}\\ 6\lambda_{13,23}&2(\lambda_{11,33}+2\lambda_{13,31})&6\lambda_{12,13}\\ 6\lambda^{*}_{12,32}&6\lambda^{*}_{12,13}&2(\lambda_{11,22}+2\lambda_{12,21})\end{bmatrix}$, (139)
+$H$ $=$ $\begin{bmatrix}6\lambda_{11,11}&2(2\lambda_{11,22}+\lambda_{12,21})&2(2\lambda_{11,33}+\lambda_{13,31})\\ 2(2\lambda_{11,22}+\lambda_{12,21})&6\lambda_{22,22}&2(2\lambda_{22,33}+\lambda_{23,32})\\ 2(2\lambda_{11,33}+\lambda_{13,31})&2(2\lambda_{22,33}+\lambda_{23,32})&6\lambda_{33,33}\end{bmatrix}$, (140)
+$I$ $=$ $\begin{bmatrix}2(\lambda_{22,33}+2\lambda_{23,32})&6\lambda^{*}_{12,32}&6\lambda_{13,23}\\ 6\lambda_{12,32}&2(\lambda_{11,22}+2\lambda_{12,21})&6\lambda_{12,13}\\ 6\lambda^{*}_{13,23}&6\lambda^{*}_{12,13}&2(\lambda_{11,33}+2\lambda_{13,31})\end{bmatrix}$, (141)
 and
 $M_{|1,1,0\rangle}=\begin{bmatrix}2(\lambda_{11,22}-\lambda_{12,21})&0&0\\ 0&2(\lambda_{11,33}-\lambda_{13,31})&0\\ 0&0&2(\lambda_{22,33}-\lambda_{23,32})\end{bmatrix}$. (142)
-V.9.2 Unitarity Bounds
-We show the results of partial wave unitarity bounds to the eigenvalues that can be computed analytically,
+V.9.2 Perturbative Unitarity Bounds
+We show the results of partial-wave unitarity bounds to the eigenvalues that can be computed analytically,
 $2|\lambda_{11,22}-\lambda_{12,21}|$ $\leq 8\pi,$ (143)
@@
 $2|\lambda_{22,33}-\lambda_{23,32}|$ $\leq 8\pi,$ (145)
-$\left|\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\pm\sqrt{\left(\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\right)^{2}-4\left(\lambda_{11,11}\lambda_{22,33}+\lambda_{11,11}\lambda_{23,32}-2|\lambda_{12,13}|^{2}\right)}\right|$ $\leq 8\pi,$ (143)
-$\left|\lambda_{11,33}+\lambda_{13,31}+\lambda_{22,22}\pm\sqrt{\left(\lambda_{11,33}+\lambda_{13,31}+\lambda_{22,22}\right)^{2}-4\left(\lambda_{11,33}\lambda_{22,22}+\lambda_{13,31}\lambda_{22,22}-2|\lambda_{12,32}|^{2}\right)}\right|$ $\leq 8\pi,$ (144)
-$\left|\lambda_{11,22}+\lambda_{12,21}+\lambda_{33,33}\pm\sqrt{\left(\lambda_{11,22}+\lambda_{12,21}+\lambda_{33,33}\right)^{2}-4\left(\lambda_{11,22}\lambda_{33,33}+\lambda_{12,21}\lambda_{33,33}-2\left|\lambda_{13,23}\right|^{2}\right)}\right|$ $\leq 8\pi.$ (145)
-Term Our Notation Notation of Bento:2017eti
+$\left|\lambda_{11,11}+\lambda_{22,33}+\lambda_{23,32}\pm\sqrt{8|\lambda_{12,13}|^{2}+\left(\lambda_{22,33}+\lambda_{23,32}-\lambda_{11,11}\right){}^{2}}\right|$ $\leq 8\pi,$ (146)
+$\left|\lambda_{22,22}+\lambda_{11,33}+\lambda_{13,31}\pm\sqrt{8|\lambda_{12,32}|^{2}+\left(\lambda_{11,33}+\lambda_{13,31}-\lambda_{22,22}\right){}^{2}}\right|$ $\leq 8\pi,$ (147)
+$\left|\lambda_{33,33}+\lambda_{11,22}+\lambda_{12,21}\pm\sqrt{8|\lambda_{13,23}|^{2}+\left(\lambda_{11,22}+\lambda_{12,21}-\lambda_{33,33}\right){}^{2}}\right|$ $\leq 8\pi.$ (148)
+Table 8: Comparison of coupling notation.
+Term Our Notation Notation of Bento et al. (2017)
 $(\Phi_{1}^{\dagger}\Phi_{1})^{2}$ $\lambda_{11,11}$ $r_{1}$
@@
 $(\Phi_{2}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{3})$ $2\lambda_{12,32}^{\ast}$ $2c_{12}^{\ast}$
-Table 8: Comparison of coupling notation.
-We compared our results with those presented in Bento:2017eti , and found them to be consistent.
-Table 8 provides a summary of the correspondence between the quartic couplings used in our study and those in Bento:2017eti .
+We compared our results with those presented in Bento et al.
+(2017), and found them to be consistent.
+Table 8 provides a summary of the correspondence between the quartic couplings used in our study and those in Bento et al.
+(2017).
 VI Overview
@@
 Such models must be subject to theoretical constraints, even before a simulation starts.
-Those constraints include boundedness from below, nonexistence of lower lying alternative vaccua, and the unitarity bounds on $2\rightarrow 2$ scattering.
-In this paper we address fully the perturbative partial-wave unitarity for the tree-level scattering matrix in models with any number of scalar doublets, neutral singlets, and/or charged singlets.
-Enforcing the correct high-energy behavior, provides bounds on the quartic couplings, freeing us from defining the exact nature of the quadratic and cubic couplings.
-In contrast, if one one wishes to turn such bounds on restrictions on masses, mixing angles and other directly observable quantities, then one must define the full theory.
-That is, one must define the quadratic and cubic couplings, a specific vacuum, and the mass matrices must be dully diagonalized.
-After this, one would strife to invert the relations, turning the unitarity bounds on the quartic couplings into restrictions on combinations of masses and mixing angles.
+Those constraints include boundedness from below, nonexistence of lower lying alternative vacua, and the perturbative unitarity bounds on $2\rightarrow 2$ scattering.
+In this paper, we address the perturbative partial-wave unitarity for the tree-level scattering matrix in models with any number of scalar doublets, neutral singlets, and/or charged singlets.
+Enforcing the correct high-energy behavior provides bounds on the quartic couplings, freeing us from defining the exact nature of the quadratic and cubic couplings.
+In contrast, if one wishes to turn such bounds into restrictions on masses, mixing angles and other directly observable quantities, then one must define the full theory.
+That is, one must define the quadratic and cubic couplings, a specific vacuum, and the mass matrices must be duly diagonalized.
+After this, one would strive to invert the relations, turning the unitarity bounds on the quartic couplings into restrictions on combinations of masses and mixing angles.
 This is possible in simple models, but increasingly more difficult as the number of fields increases.
 In contrast, although the matrices get larger, our limits on quartic couplings are always applicable, at least numerically.
-We classify the states by the conserved quantum numbers $Q$, $Y$, and $T$ and show that, once one restricts oneself to the minimal set of states providing all inequivalent bounds, the quantum number $Q$ is redundant.
-We also discuss examples where the existence of extra symmetries allows for the inclusion of further quantum numbers in the basis, thus greatly simplifying the matrices.
+We classify the states by the conserved quantum numbers $Q$, $Y$, and $T$, and show that, once one restricts oneself to the minimal set of states providing all inequivalent bounds, the quantum number $Q$ is redundant.
+We also discuss examples where the existence of extra symmetries allows for the inclusion of further quantum numbers in the basis, thus greatly simplifying the scattering matrices.
 We introduce the Mathematica notebook BounDS that automatically calculates the quartic part of the potential and the scattering matrices and their eigenvalues, for any model with any symmetries (discrete or continuous, Abelian or non-Abelian).
@@
 For this reason, in building $2\to 2$ scattering matrices, Ref.
-Bento:2017eti labeled two-particle states by $|Q,Y\rangle$.
+Bento et al.
+(2017) labeled two-particle states by $|Q,Y\rangle$.
 For models like the one defined in Section II, we list in Table.
 9 all possible two-particle states.
+Table 9: Basis of two-particle states labeled by $|Q,Y\rangle$. This table includes all states, some of which provide redundant information.
 $|Q,Y\rangle$ State Conditions Dimensionality
@@
 $|0,0\rangle$ $\left\{\phi_{i}^{+}\phi_{j}^{-},\phi_{i}^{0}\phi_{j}^{0*},\ \varphi_{i}^{+}\varphi_{j}^{-},\ \chi_{i}\chi_{j}\right\}$ $\{\,\textrm{---}\,,\,\textrm{---}\,,\,\textrm{---}\,,\,i\leq j\}$ $2n_{D}^{2}+n_{c}^{2}+n_{n}(n_{n}+1)/2$
-Table 9: Basis of two-particle states labeled by $|Q,Y\rangle$. This table includes all states, some of which provide redundant information.
 However, not all scattering amplitudes are independent. Notice that
-$\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\delta_{ca,db},$ (146)
-$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\phi^{0}_{b}\to\phi^{0}_{c}\phi^{0}_{d}\right]=2\lambda_{ca,db}+2\lambda_{da,cb},$ (147)
+$\mathcal{M}\left[\phi^{+}_{a}\varphi^{+}_{b}\to\phi^{+}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\varphi^{+}_{b}\to\phi^{0}_{c}\varphi^{+}_{d}\right]=\delta_{ca,db},$ (149)
+$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\phi^{0}_{b}\to\phi^{0}_{c}\phi^{0}_{d}\right]=2\left(\lambda_{ca,db}+\lambda_{da,cb}\right),$ (150)
 $\mathcal{M}\left[\phi^{+}_{a}\chi_{b}\to\phi^{+}_{c}\chi_{d}\right]=\mathcal{M}\left[\phi^{0}_{a}\chi_{b}\to\phi^{0}_{c}\chi_{d}\right]=2\gamma_{ca,bd},$ (151)
@@
 $\mathcal{M}\left[\phi^{0*}_{a}\varphi^{+}_{b}\to\phi^{0*}_{c}\varphi^{+}_{d}\right]=\mathcal{M}\left[\phi^{-}_{a}\varphi^{+}_{b}\to\phi^{-}_{c}\varphi^{+}_{d}\right]=\delta_{ac,db}.$ (153)
-Therefore, perturbative unitarity bounds obtained from scatterings involving $|1,\frac{3}{2}\rangle$, $|0,1\rangle$ , and $|0,\frac{1}{2}\rangle$ are redundant because they are identical to those derived from $|2,\frac{3}{2}\rangle$, $|2,1\rangle$ , and $|1,\frac{1}{2}\rangle$ , respectivelly.
-In any scattering involving $SU(2)$ doublets, total isospin $T$ must also be conserved, For this reason, two-particle states should be further labeled by $|Q,Y,T\rangle$, with $T=0,1$ in this model.
+Therefore, perturbative unitarity bounds obtained from scatterings involving $|1,\frac{3}{2}\rangle$, $|0,1\rangle$ , and $|0,\frac{1}{2}\rangle$ are redundant because they are identical to those derived from $|2,\frac{3}{2}\rangle$, $|2,1\rangle$ , and $|1,\frac{1}{2}\rangle$ , respectively.
+In any scattering involving $SU(2)$ doublets, total isospin $T$ must also be conserved, For this reason, two-particle states should be further labeled by $|Q,Y,T\rangle$, with $T=0,1$ in this class of models.
 Using Clebsch-Gordan coefficients, we can split the two-particle state $\phi^{+}_{i}\phi^{0}_{j}$ in $|1,1\rangle$ into
@@
 These redefinitions of states lead, once again, to redundant scattering matrices because
-$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{(a}\,\phi^{0}_{b)}\to\phi^{+}_{(c}\,\phi^{0}_{d)}\right]=2\lambda_{ca,db}+2\lambda_{da,cb},$ (155)
+$\mathcal{M}\left[\phi^{+}_{a}\phi^{+}_{b}\to\phi^{+}_{c}\phi^{+}_{d}\right]=\mathcal{M}\left[\phi^{+}_{(a}\,\phi^{0}_{b)}\to\phi^{+}_{(c}\,\phi^{0}_{d)}\right]=2\left(\lambda_{ca,db}+\lambda_{da,cb}\right),$ (158)
 $\mathcal{M}\left[\phi^{+}_{a}\phi^{0*}_{b}\to\phi^{+}_{c}\phi^{0*}_{d}\right]=\mathcal{M}\left[\overline{\Phi_{a}\Phi_{b}^{*}}\to\overline{\Phi_{c}\Phi_{d}^{*}}\right]=2\lambda_{ca,bd}.$ (159)
-We therefore conclude that it is sufficient to apply the partial wave unitarity bounds to the scattering matrices built out of the states listed in Table 1, i.e. the states
+We therefore conclude that it is sufficient to apply the partial-wave unitarity bounds to the scattering matrices built out of the states listed in Table 1, i.e. the states
 $\left|Q,Y,T\right\rangle\quad=\quad\left|2,2,0\right\rangle,\ \left|2,\frac{3}{2},\frac{1}{2}\right\rangle,\ \left|2,1,1\right\rangle,\ \left|1,1,0\right\rangle,\ \left|1,\frac{1}{2},\frac{1}{2}\right\rangle,\ \left|1,0,1\right\rangle,\ \left|0,0,0\right\rangle.$ (160)
@@
 $|0,0\rangle$ $\left\{\phi^{+}_{1}\phi^{-}_{1},\ \phi^{0}_{1}\phi^{0*}_{1}\right\}$ 2
-Table 10: Basis of two-particle states labelled by $|Q,Y\rangle$.
+Table 10: Basis of two-particle states labeled by $|Q,Y\rangle$.
 The corresponding scattering matrices are
@@
 $|0,1,1\rangle$ $\phi^{0}_{1}\phi^{0}_{1}$ 1
-Table 11: Basis of two-particle states labelled by $|Q,Y,T\rangle$.
+Table 11: Basis of two-particle states labeled by $|Q,Y,T\rangle$.
 Let transformation from the old $(Q,Y)$ basis to the new $(Q,Y,T)$ basis be given by
@@
 Appendix C BounDS
-We have developed BounDS, a Mathematica notebook, that automates the process of writing the quartic potential and deriving partial wave unitarity bounds for models within the class of models defined in Section II.
-Notice that ScannerS Coimbra:2013qq ; Muhlleitner:2020wwk has a tool to calculate the scattering matrices, but the user must introduce the correct potential.
+We have developed BounDS, a Mathematica notebook, that automates the process of writing the quartic potential and deriving partial-wave unitarity bounds for models within the class of models defined in Section II.
+Notice that ScannerS Coimbra et al.
+(2013); Mühlleitner et al.
+(2022) has a tool to calculate the scattering matrices, but the user must introduce the correct potential.
 Here, the user just introduces the number of fields and their symmetries; our program calculates the correct quartic potential and then the relevant scattering matrices.
@@
 The list Assume contains information about group theoretical parameters, such as the space they are defined in and relationships between them.
-For these parameters, the user must use Latin characters.
 Furthermore, the function Conjugate[] may be used for symmetries that require conjugation of a field.
 BounDS has been successfully tested with a wide range of symmetry groups, including both discrete and continuous, Abelian and non-Abelian cases.
+The list Sym[] is read sequentially, and some symmetries may take longer to resolve than others.
+For improved time performance, we recommend that the user declare Abelian symmetries before imposing possible non-Abelian symmetries.
 After evaluating the cells in step 1, the user can simply run all the cells in step 2. Under the hood, this module does the following operations:
@@
 By calling Potential4, the user can output the quartic part of the most general scalar potential allowed by the symmetries of the model.
-Anywhere in the code, the Mathematica function TeXForm can be called to get the source code:
+Anywhere in the code, the native Mathematica function TeXForm can be called to get the source code:
 Potential4 //TeXForm
@@
 In a process involving complex scalars that are singlets under the SM gauge group, there are CP-even and CP-odd components that scatter independently.
-Furthermore, if there are additional flavour symmetries, each two-particle state must also be labelled by its corresponding charge $S_{\alpha}$, following the idea of Ginzburg:2005dt .
+Furthermore, if there are additional flavour symmetries, each two-particle state must also be labeled by its corresponding charge $S_{\alpha}$, following the idea of Ginzburg and Ivanov (2005).
 Therefore, we can label all states by,
@@
 $|0,0,0,-1\rangle$ $\{\Phi_{1}\Phi_{2}^{*},\,\Phi_{2}\Phi_{1}^{*}\}$ 2
-Table 12: Basis of two-particle states labelled by $|Q,Y,T,\mathbb{Z}_{2}\rangle$.
+Table 12: Basis of two-particle states labeled by $|Q,Y,T,\mathbb{Z}_{2}\rangle$.
 The scattering matrices are thus
@@
 Using the substitutions in Table 4 we recover the well-known unitarity bounds for the $\mathbb{Z}_{2}$-symmetric 2HDM, shown in Eqs. (65)–(70).
-We find that, by labelling states with flavour symmetries in addition to charge, hypercharge, and total isospin, the scattering matrices simplify further and reduce in dimensionality compared to those in Section V.3.1.
+We find that, by labeling states with flavour symmetries in addition to charge, hypercharge, and total isospin, the scattering matrices simplify further and reduce in dimensionality compared to those in Section V.3.1.
 This reduction greatly facilitates the subsequent calculations, as shown in this appendix.